### v1

# CO3: Contrasting Concepts Compose Better

Abstract
We propose to improve multi‑concept prompt fidelity in text‑to‑image diffusion models. We begin with common failure cases—prompts like “a cat and a dog” that sometimes yields images where one concept is missing, faint, or colliding awkwardly with another. We hypothesize that this happens when the diffusion model drifts into mixed modes that over‑emphasize a single concept it learned strongly during training. Instead of re-training, we introduce a corrective sampling strategy that steers away from regions where the joint prompt behavior overlaps too strongly with any single concept in the prompt. The goal is to steer towards “pure” joint modes where all concepts can coexist with balanced visual presence.. We further show that existing multi‑concept guidance schemes can operate in unstable weight regimes that amplify imbalance; we characterize favorable regions and adapt sampling to remain within them. Our approach, CO3, is plug‑and‑play, requires no model tuning, and complements standard classifier‑free guidance. Experiments on diverse multi-concept prompts indicate improvements in concept coverage, balance and robustness, with fewer dropped or distorted concepts compared to standard baselines and prior compositional methods. Results suggest that lightweight corrective guidance can substantially mitigate brittle semantic alignment behavior in modern diffusion systems.
1 Introduction
Recent diffusion models (Ho et al., 2020; Ramesh et al., 2022a; Rombach et al., 2022) have ushered significant breakthroughs in Text-to-Image (T2I) synthesis, producing high-fidelity images from textual descriptions. However, ensuring the generated images faithfully adhere to the prompt, a challenge known as semantic alignment (Chefer et al., 2023; taihang Hu et al., 2024; Liu et al., 2023), remains a problem. Concretely, for a given prompt $C$, T2I models like StableDiffusion (Rombach et al., 2022) sample from the modes (or high probability regions) of the learned distribution, $p(x|C)$. While such models can produce high resolution images in general, every so often, the results are surprisingly misaligned even for very simple prompts containing few concepts, e.g., CC=“a cat and a dog”. Diagnosing exactly why this behavior emerges periodically is difficult. It is conceivable that the complex training process in high dimensions, especially in conjunction with text embeddings, creates some problematic modes in $p(x|C)$.
We hypothesize that problematic modes in $p(x|C)$ arise when they overlap with modes of individual concept distribution $p(x|c_{i})$. Such an overlap biases the generation toward a single concept $c_{i}$, reducing the prominence of others. For instance, across images of $c_{1}=\text{``cat''}$ in the training dataset, a few may have an inconspicuous or partial $c_{2}=\text{``dog''}$ in the background. This image may still fall under the mode of $p(x|C)$. We attribute this to training instabilities and relatively less coverage of multi-concept prompts $C$, which cause the model to assign high probability even to weakly conforming images. Said differently, an image of a big cat and an inconspicuous dog can get assigned high probabilities under $p(x|C)$, causing semantic misalignment.
Preventing such problematic modes warrants strict and specialized training paradigms; a difficult task for such large models. However, “curing” them after their occurrence is a more viable approach Assuming our hypothesis is true, we propose a cure for problematic modes. Our intuitive idea is to go away from problematic modes and move towards modes under which none of the individual concepts are strong. To realize this, we propose to design a corrector that generates samples from the following distribution:
$\tilde{p}(x\mid C)\;\propto\;\frac{p(x\mid C)}{\prod_{i}p(x\mid c_{i})}.$ (1)
Figure 1 illustrates the intuition behind our proposal. Our corrector distribution $\tilde{p}(x|C)$ assigns low probability to regions where $p(x|C)$ overlaps with individual $p(x|c_{i})$; we deem them as degenerate modes dominated by a single concept. By suppressing these overlaps, the corrector emphasizes pure $p(x|C)$ modes where all concepts coexist without one overwhelming the others. From a probabilistic perspective, this acts as a corrective factor: while $p(x|C)$ may assign high probability to weakly conforming images due to training noise or limited multi-concept data, dividing by the marginals removes this bias and sharpens the distribution toward genuine multi-concept samples. As a result, the modes we target are more semantically aligned and less prone to concept suppression or distortion.
Correction sampling from $\tilde{p}(x|C)$ can be achieved by composing scores from constituent component distributions $\nabla_{x}\log p(x|C),\{\nabla_{x}\log p(x|c_{i})\}_{i}$ (Liu et al., 2022). While there are many ways to compose, we analyse and show that composition through weighted sum of Tweedie-means—in the Tweedie denoised space—offers a more general framework that subsumes existing approaches.
In particular, we show that two classes of correction sampling—noise-resampling and latent correction (Bao et al., 2024; Rassin et al., 2023; taihang Hu et al., 2024; Kwon & Ye, 2025)—become special cases of Tweedie-mean composition under different weighting schemes. This allows us to design a hybrid composition framework that serves the purpose of resampling at time $T$, and then toggles to latent correction at later steps. By latent, we mean that the correction steps are accomplished in between each DDIM time step, allowing CO3 to be a plug-and-play, model-agnostic and gradient-free approach for T2I models. Comparison of CO3 against SOTA baselines shows stronger semantic alignment to prompts (measured using multiple metrics), giving empirical evidence of our hypothesis.
Figure 1: The figure illustrates our hypothesis on mode overlap using a simple 2D toy example. (a) Two modes of the distribution $p_{t}(x|\textit{"a cat and a dog"})$ (in green contour) has significant overlap with the modes of the individual concept distributions $p_{t}(x|\textit{"a cat"})$ (in red contour) and $p_{t}(x|\textit{"a dog"})$ (in orange contour). (b) The proposed corrector distribution $p_{t}(x|\texttt{"a cat and a dog"})/(p_{t}(x|\textit{"a cat"})p_{t}(x|\textit{"a dog"}))$ suppresses these overlaps, steering the generation away from problematic modes. The arrows indicate the denoising directions.
2 Background
$\blacksquare$ Conditional generation using Classifier-Free Guidance (CFG). In diffusion-based Text-to-Image (T2I) generation (Rombach et al., 2022; Saharia et al., 2022a; Ramesh et al., 2022b), given the noisy latent $x_{t}$ at timestep $t$, a denoised estimate can be derived using Tweedie’s formula:
$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(x_{t},c,t)}{\sqrt{\bar{\alpha}_{t}}},$ (2)
where $\epsilon_{\theta}$ denotes the predicted noise conditioned on the text prompt $c$, and $\bar{\alpha}_{t}$ is the cumulative product of the noising schedule. This step corresponds to the denoising stage, recovering an estimate of the clean signal $x_{0}$. In the DDIM sampler (Song et al., 2021), under the noise-free condition, the subsequent step deterministically evolves $\hat{x}_{0}$ to $x_{t-1}$ without introducing additional stochasticity:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\,\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\,\epsilon_{\theta}(x_{t},c,t).$ (3)
Here, the same predicted noise $\epsilon_{\theta}$ is reused, eliminating the renoisification step present in stochastic samplers such as DDPM(Ho et al., 2020). This perspective highlights how the Tweedie decomposition denoising via $\hat{x}_{0}$, followed by deterministic reconstruction of $x_{t-1}$, is naturally aligned with DDIM sampling.
In practice, most T2I models adopt classifier-free guidance (CFG) (Ho & Salimans, 2022), where predictions from both the conditional and unconditional models are combined:
$\epsilon_{t}^{\lambda,c}\;=\;\lambda\ \epsilon_{\theta}(x_{t},c,t)\;+\;(1-\lambda)\,\epsilon_{\theta}(x_{t},\varnothing,t)$ (4)
where $\lambda>1$ controls the guidance strength. $\epsilon_{t}^{\lambda,c}$ is the composed noise prediction under CFG. Then the denoising and DDIM steps proceed as before, but using $\epsilon_{t}^{\lambda,c}$ in place of $\epsilon_{\theta}(x_{t},c,t)$.
Tweedie View with CFG.
Substituting $\epsilon_{t}^{\lambda,c}$ into the Tweedie denoising and DDIM update yields:
$\hat{x}_{0}^{\lambda,c}$ $=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}^{\lambda,c}}{\sqrt{\bar{\alpha}_{t}}}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}],$ (5)
$x_{t-1}$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\,\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}]+\sqrt{1-\bar{\alpha}_{t-1}}\,\epsilon_{t}^{\lambda,c}$ (6)
where $\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}]:=x_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}^{\lambda,c}$ is the Tweedie mean from the CFG noise at $t$. Thus, CFG in the Tweedie framework can be interpreted as modifying the denoised estimate $\hat{x}_{0}$ to $\hat{x}_{0}^{\lambda,c}$ before the deterministic step to $x_{t-1}$ (Chung et al., 2024; Kwon & Ye, 2023).
$\blacksquare$ Correction-based approaches for conditional generation. A number of recent works have addressed the challenge of compositional text-to-image generation using correction-based approaches (Chefer et al., 2023; Liu et al., 2023; Bao et al., 2024; Rassin et al., 2023; taihang Hu et al., 2024). During sampling with classifier-free guidance, these methods iteratively correct the latent variable by applying gradient updates of the form
$x^{k+1}_{t}=x^{k}_{t}-s\,\nabla_{x_{t}}\mathcal{L}(x_{t},c),\quad k=1,2,\ldots,M-1.$ (7)
and then use the final refined latent $x^{M}_{t}$ to predict the next DDIM step $x_{t-1}$, using Eqs. 5 and 6. Here $\mathcal{L}(x_{t},c)$ is a task-specific loss function that enforces better alignment with the target condition $c$, and $s$ is a step size. This iterative update can be interpreted as correcting the diffusion guidance process using a corrector distribution of the form:
$\tilde{p}_{t}(x_{t},c)\propto\exp\bigl(-\mathcal{L}(x_{t},c)\bigr),$ (8)
which refines the generative process at each timestep. In the special case where the step size $s=\sigma^{2}_{t}$ $1-\bar{\alpha}_{t}$ (the noise scale at timestep $t$), the update rule becomes equivalent to iterative Tweedie-mean correction (see Eq. 5 above), where $x_{t}^{k+1}$ is reinterpreted as the next estimated Tweedie mean given $x_{t}^{k}$ for the distribution $\tilde{p}_{t}(x_{t},c)$.
$\blacksquare$ Composable-diffusion. Generating samples that satisfy multiple conditions $\{c_{1},\ldots,c_{K}\}$ can be formulated as sampling from the joint distribution
$\tilde{p}_{0}(x_{0}\mid c_{1},\ldots,c_{K})\;\propto\;\prod_{k=1}^{K}p_{0}(x_{0}\mid c_{k}).$ (9)
To achieve this, Liu et al. (2022) proposed composable diffusion, which directly composes the output scores (predicted noises) from different conditional diffusion models using CFG during sampling.
Specifically, in the text-to-image setting, if the prompt $C$ can be decomposed into constituent concepts $\{c_{1},c_{2},\ldots,c_{K}\}$, the outputs of the corresponding diffusion chains can be combined as
$\tilde{\epsilon}^{\lambda,C}_{t}\;=\;\epsilon_{t}^{\phi}+\lambda_{1}\bigl(\epsilon_{t}^{c_{1}}-\epsilon_{t}^{\phi}\bigr)+\lambda_{2}\bigl(\epsilon_{t}^{c_{2}}-\epsilon_{t}^{\phi}\bigr)+\dots+\lambda_{K}\bigl(\epsilon_{t}^{c_{K}}-\epsilon_{t}^{\phi}\bigr)$ (10)
where $\epsilon_{t}^{\phi}$ denotes the unconditional score, and $\lambda_{k}$ controls the guidance strength for concept $c_{k}$. The next sample is then predicted via the Tweedie formulation:
$x_{t-1}=\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\,\hat{x}_{\text{tweedie}}\!\left[\tilde{\epsilon}^{\lambda,C}_{t}\right]\;+\;\sqrt{1-\bar{\alpha}_{t-1}}\,\tilde{\epsilon}^{\lambda,C}_{t}.$ (11)
Although this approach is model-agnostic and conceptually simple, its performance is often poor, since the above linear composition of scores is incorrect or doesn’t correspond to the score of the diffusion forward distribution $\tilde{p}_{t}(x_{t}\mid c_{1},\ldots,c_{K})$ at any timestep $t>0$ (Du et al., 2024).
In summary, both the correction-based approach and the composable diffusion idea can be interpreted as different ways of approximating Tweedie-means $\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}]$ at time $t$.
3 CO3: Contrasting Concepts to Compose
We aim to combine the strengths of correction-based approaches and composable diffusion. On one hand, correction-based methods are powerful and explicitly improve compositional alignment, but they are subject to the complex gradient of the base model. On the other hand, composable diffusion is fully model-agnostic, but suffers from poor performance because linear score composition is not consistent with the diffusion forward process.
To take advantage of both, assume that prompt $C$ can be decomposed into constituent concepts $\{c_{1},c_{2},\ldots,c_{K}\}$. We propose an explicit Concept Contrasting corrector distribution based on our hypothesis on mode overlap discussed in Sec. 1. Specifically, we define the corrector distribution at each timestep $t$ as:
$\tilde{p}_{t}(x_{t},C)\propto\frac{p_{t}(x_{t}\mid C)^{w_{0}}}{\prod_{k=1}^{K}p_{t}(x_{t}\mid c_{k})^{w_{k}}},$ (12)
where $\{w_{0},w_{1},\ldots,w_{K}\}$ are composition weights. As discussed in Sec. 1, this corrector steers the generation toward regions where the distribution $p_{t}(x_{t}\mid C)$ is high, while simultaneously avoiding regions of overlap with the individual concept distributions $\{p_{t}(x_{t}\mid c_{k})\}_{k}$. Observe that this encourages the model to generate samples that satisfy all concepts in $C$ without over-emphasizing any single concept. We present the CO3 corrector pseudo code in Algorithm 1.
To sample from the unnormalized probability distribution in Eq. 12, an well-known approach is to compose the concept distributions $\{p_{t}(x_{t}\mid c_{k})\}_{k}$ in the space of score functions (Liu et al., 2022; Du et al., 2024). We break-away from this approach and compose the distribution in the Tweedie mean space; we show how this offers a more general framework for composition.
In Tweedie-denoised space we define composition as:
$\tilde{x}_{tweedie}=w_{0}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,C}]\;+\;w_{1}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{1}}]\;+\;\dots\;+\;w_{K}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{K}}],$ (13)
Algorithm 1 DDIM with CO3 Corrector
[…]
2:for $t=T$ down to 11 do
3:  if $t>T_{c}$ then
4:   $x_{t}^{(2)}\leftarrow\text{{CO3}}(x_{t},C,c_{1},\dots,c_{k})$
5:    $\triangleright$ Proposed correction
6:   $x_{t}\leftarrow x_{t}^{(2)}$
7:  end if
8:  $x_{t}\leftarrow\text{DDIM\_step}(x_{t})$ $\triangleright$ Inverse DDIM
9:end for
where $w_{0}>0\text{ and }w_{1},\dots,w_{K}<0$ are concept weights. Note: $\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{k}}]$ is the Tweedie mean corresponding to the CFG composed noise prediction $\epsilon_{t}^{\lambda,c_{k}}$ for concept $c_{k}$ with guidance weight $\lambda$.
How should composition weights be chosen? We analyze the effect of different weight assignments to Eq. 13, in particular, how the constraint on the concept weights $w_{i}$} influences both the theoretical interpretation and the empirical behavior of the compositional Tweedie mean.
A valid compositional Tweedie mean $\tilde{x}_{tweedie}$ in Eq. 13 must be one that can be expressed in the definition given in equation 5 as DDIM goes through a series of Tweedie means for image generation as described in section 2. Lemma 1 shows that this is guaranteed only when the weights satisfy the normalization condition $\sum_{i}w_{i}=1$.
Lemma 1.
Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$. Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$. Then,
  1. a)
    CO3-corrector: $\tilde{\hat{x}}_{tweedie}$ can be expressed in the form of a tweedie-mean at time $t$, i.e. $\tilde{\hat{x}}_{tweedie}=x-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ if and only if $\sum_{k}w_{k}=1$. Here $\tilde{\lambda}=\lambda$ and CFG composed noise $\tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$.
  2. b)
    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
[…]
To further characterize the behavior of the composition under the two conditions $\sum_{k}w_{k}=1$ and $\sum_{k}w_{k}=0$, we evaluate the performance of different weighting strategies across timesteps $t$ of the diffusion chain. Figure 2 reveals complementary dynamics between the resampler and the corrector. Specifically, the resampler ($\sum_{k}w_{k}=0$) is effective at high $t$ values near $T$, but its effectiveness diminishes as the denoising process progresses toward lower $t$. Although resampling is not theoretically valid for $t<T$, we observe practical improvements until approximately $t\approx 0.9T$. In contrast, the corrector ($\sum_{k}w_{k}=1$) exhibits gradual improvement with increasing timesteps and saturates around $t\approx 0.75T$. These findings invite a hybrid strategy into CO3: apply resampling during early timesteps ($t>T_{R}$), followed by correction until a threshold ($T_{C}<t<T_{R}$), beyond which further gains are marginal.
The pseudocode for the zero-sum-weight resampler (CO3-resampler) and the unit-sum-weight (CO3-corrector) corrector is provided in Algorithm 2 and Algorithm 3, respectively. Importantly, the same diffusion denoiser model can be reused to compute the individual concept scores under different input conditions $c_{k}$}, thereby eliminating the need for any costly backward passes to obtain gradients. Furthermore, the framework operates without reliance on model-specific architectural details, making it fully model-agnostic and gradient-free.
Figure 2: Characterization of Resampler and Corrector steps. Resampling is more powerful at high $t$ while the Corrector improves slowly with more timesteps and saturates.
Closeness-Aware Concept Weight Modulation: To make the CO3 corrector more adaptive, we anchor the weight $w_{0}$ ($w_{0}$=1.0 in Algo. 2 and $w_{0}=2.0$ in Algo. 3), and assign weights to each concept based on how close the current noise prediction $\epsilon^{C}$ is, to the noise corresponding to that concept, $\epsilon^{c_{k}}$. Intuitively, if a sample looks closer to concept $c_{k}$ than to all others, we want to penalize $c_{k}$ more strongly (i.e., give it a larger negative weight), while reducing the strength of the other concepts. This encourages the sampler to move away from the nearest mode, preventing collapse toward one dominant concept. Formally, let $d_{k}=\lVert\epsilon^{C}-\epsilon^{c_{k}}\rVert$ denote the distance between the current sample and the mode of concept $c_{k}$. We convert distances $\{d_{k}\}_{k=1}^{K}$ into affinity scores using an exponential kernel:
$a_{k}=\exp(-\beta\,d_{k}),\qquad\beta>0$ (14)
so that concepts closer to the current sample (smaller $d_{k}$) receive higher affinity. These affinity scores are then normalized to define the weights:
$w_{k}=-\frac{a_{k}}{\sum_{j=1}^{K}a_{j}},$ (15)
which ensures each weight is negative and the total sum satisfies $\sum_{k=1}^{K}w_{k}=-1$. As a result, concepts that are closer to the current sample receive stronger negative weights, while farther concepts are down-weighted in proportion to their distance. This weight modulation scheme samples from the whole hyper-plane $\sum_{k=0}^{K}w_{k}=1$ (for CO3-correction) or $\sum_{k=0}^{K}w_{k}=0$ (for CO3-resampling) instead of a fixed weights for the entire course of generation.
Algorithm 2 CO3-resampler ($\sum w_{k}=0$)
[…]
2:for $p=1$ to $P$ do
3:  ${\epsilon}^{c_{0}}=\psi(x_{t},C)$, ${\epsilon}^{c_{k}}=\psi(x_{t},c_{k})$,
4:      $k=1,\dots,K$
5:  $\hat{x}_{tweedie}[\epsilon^{\lambda,c_{k}}_{t}]=x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}^{\lambda,c_{k}}$
6:  $\tilde{\hat{x}}_{tweedie}=-\lambda\sqrt{1-\bar{\alpha}_{t}}\sum_{k=0}^{K}w_{k}{\epsilon}^{c_{k}}$
7:   $\triangleright$ Projection onto noise space
8:  $x_{t}^{(2)}=\tilde{\hat{x}}_{tweedie}+\sqrt{1-\bar{\alpha}_{t}}\ \epsilon_{t}^{\phi}$
9:   $\triangleright$ Uncond. manifold correction
10:  $x_{t}\leftarrow x_{t}^{(2)}$
11:end for
12:return $x_{t}$
Algorithm 3 CO3-corrector ($\sum w_{k}=1$)
[…]
2:for $p=1$ to $P$ do
3:  ${\epsilon}^{c_{0}}=\psi(x_{t},C)$, ${\epsilon}^{c_{k}}=\psi(x_{t},c_{k}),$
4:      $k=1,\dots,K$
5:  $\tilde{\hat{x}}_{tweedie}=\sum_{k=0}^{K}w_{k}\hat{x}_{tweedie}[{\epsilon}^{\lambda,c_{k}}_{t}]$
6:   $\triangleright$ Tweedie mean composition
7:  $r=\frac{\lVert\hat{x}_{tweedie}^{c_{0}}\rVert_{2}}{\lVert\tilde{\hat{x}}_{tweedie}\rVert_{2}}$ $\triangleright$ mean normalization
8:  $\tilde{\hat{x}}_{tweedie}\leftarrow\tilde{\hat{x}}_{tweedie}*r$
9:  $x_{t}^{(2)}=\tilde{\hat{x}}_{tweedie}+\sqrt{1-\bar{\alpha}_{t}}\ \epsilon_{t}^{\phi}$
10:  $x_{t}\leftarrow x_{t}^{(2)}$
11:end for
12:return $x_{t}$
4 Experiments
4.1 Experimental Setup
Implementation details: We implement our method on SDXL (Podell et al., 2023) using a 50-step DDIM sampler with guidance scale $\lambda=5.0$. CO3 is applied during the first $20\%$ of denoising steps ($T_{c}=0.8T$), with the initial three steps reserved for resampling. For weight modulation, we set $\beta=0.8$. All experiments are performed on a single NVIDIA A6000 GPU (48GB). Additional hyperparameter details are provided in the Appendix section A.3.
Evaluation benchmark and metrics: We evaluate our method on the benchmark dataset from Chefer et al. (2023), which includes three categories of prompts: Animal–Animal, Animal–Object, and Object–Object. While the benchmark is limited to these cases, CO3 naturally extends to more complex prompts involving arbitrary numbers of concepts. For evaluation, we adopt BLIP-VQA (Huang et al., 2023) and ImageReward (Xu et al., 2023), both widely used in recent works (taihang Hu et al., 2024; Hu et al., 2024) to measure the faithfulness of generated images to input prompts. Notably, ImageReward is a learned reward model trained on human preferences, capturing aspects such as prompt adherence, aesthetics, and overall quality. For both metrics, higher scores indicate better alignment.
Comparison methods: We compare CO3 against state-of-the-art approaches, including optimization-based correction methods (Attend-and-Excite (Chefer et al., 2023), SynGen (Rassin et al., 2023), Divide-and-Bind (Liu et al., 2023), ToME (taihang Hu et al., 2024)), the composable generation method Composable Diffusion (Liu et al., 2022), and the noise-optimization method InitNO (Guo et al., 2024). All optimization-based approaches are architecture-dependent and require gradient computations at inference, with InitNO being model-agnostic but still incurring costly backpropagation. In contrast, only Composable Diffusion and our proposed CO3 are purely sampling-based, model-agnostic, and completely gradient-free.
4.2 Experimental Results
Properties BLIP-VQA ImageReward
training-free gradient-free model-agnostic animals animals-objects objects animals animals-objects objects
SD1.5(Rombach et al., 2022) - 0.3239 0.5958 0.2730 -0.2733 0.4262 -0.5521
Attend-excite(Chefer et al., 2023) 0.6980 0.7865 0.5155 0.8244 1.2380 0.8741
SynGenRassin et al. (2023) 0.3348 0.7689 0.6595 -0.2445 0.9838 0.7315
Divide & Bind 0.7201 0.8399 0.5887 0.8499 1.2516 0.8134
InitNO(Guo et al., 2024) 0.7264 0.7998 0.5406 1.0082 1.3927 1.1383
SDXLPodell et al. (2023) - 0.6950 0.8654 0.4926 0.7820 1.5574 0.6789
SynGen(SDXL)Rassin et al. (2023) 0.6816 0.8578 0.4652 0.6998 1.5622 0.6441
ToMe(taihang Hu et al., 2024) 0.6257 0.8808 0.6440 0.3895 1.5736 1.0118
Compose-diffusion(Liu et al., 2022) 0.2846 0.5656 0.4529 -1.1399 -0.2068 -0.0955
CO3(Ours) 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158
Table 1: Quantitative comparison of different methods on compositional generation tasks. We evaluate the generated images using two metrics: BLIP-VQA and ImageReward. Top performing model is highlighted in Black and 2nd best in Blue. Higher the score the better.
Quantitative Comparison:
Table 1 demonstrates the extent to which CO3, despite being model-agnostic and optimization-free, competes or outperforms optimization-based and model-agnostic baselines on ImageReward. Gains are substantive in the Animal–Animal and Animal–Object categories. On BLIP-VQA, CO3 delivers the best performance in these two categories and improves over the base SDXL on the Object category, which contains prompts of the form: “a [color1] [object1] and a [color2] [object2].”
Unlike methods such as ToMe taihang Hu et al. (2024), SynGen (Rassin et al., 2023), Attend-and-Excite (Chefer et al., 2023), and Divide-and-Bind Liu et al. (2023), which rely on model-specific architectures or explicit attention-based binding losses, CO3 is architecture-independent, sampling-based, and free of explicit subject–attribute binding. Interestingly, by simply targeting “pure” modes under $p(x\mid C)$, CO3 discovers layouts that better capture multi-object relations (Animals–Animals, Animal–Objects) and finds high-probability regions that preserve subject–attribute bindings (Objects). This performance gain empirically validates our hypothesis on mode overlap and also highlights CO3’s effectiveness in correcting the compositional limitations of SDXL.
Qualitative Comparison: Figure 3 compares CO3 with several baselines. Columns 1–2 show results from Divide-Bind (Liu et al., 2023) and Attend-and-Excite (Chefer et al., 2023) (based on SDv1.5 (Rombach et al., 2022)), while the remaining columns compare against Composable Diffusion (Liu et al., 2022), SynGen (Rassin et al., 2023), and ToMe (taihang Hu et al., 2024). We illustrate common issues in compositional alignment—concept missing, attribute mixing, and object binding—across all categories. Divide-Bind and Attend-and-Excite enhance the attention layer of the base model, but often yield cropped or incomplete concepts (rows 1, 3). Composable Diffusion suffers from missing or merged concepts (rows 2, 3). SynGen and ToMe perform well on Object prompts of the form “a [color1] [object1] and a [color2] [object2],” but still exhibit overlapping or mixed concepts (rows 2, 3; columns 5, 6). In contrast, CO3 achieves stronger object binding in the Animals and Animal–Objects categories while also improving results on Objects.
Figure 3: Qualitative comparison of different methods on simpler prompts.
Figure 4 extends qualitative results to complex prompts from the Complex category of Huang et al. (2023), comparing CO3 with the top-performing SDXL models (Table 1). The first row shows prompts involving spatial relations (e.g., “on,” “to the left of”), while the second row depicts more descriptive prompts with multiple interacting concepts. SynGen and ToMe successfully bind concepts to their subjects but often miss concepts or relations. By mitigating dominant-concept bias, CO3 remains effective even for complex prompts with multiple relations.
Figure 4: Qualitative comparison of CO3 with competing methods on complex prompts.
Ablations Studies are preformed on the key components of CO3, starting from the base SDXL model (Table 2). Adding resampling in the first three DDIM steps substantially improves average performance, particularly for the Animals and Animal–Objects categories. Since image layout is largely determined in the early timesteps, CO3-resampler iteratively explores better initial noise configurations that align with the prompt. Incorporating CO3-corrector over the next seven steps further improves prompts requiring fine-grained details such as color or texture, boosting performance on the Objects category in both ImageReward and BLIP-VQA. Additional gains come from the Closeness-Aware Concept Weight Modulation strategy, which consistently benefits all categories. We hypothesize that as the diffusion trajectory evolves, the latent state often drifts closer to one concept mode $c_{i}$; the modulation mechanism assigns $c_{i}$ a more negative weight, repelling the latent and mitigating concept dominance (Tunanyan et al., 2023) commonly observed in stable-diffusion.
BLIP-VQA ImageReward
Methods animals animals-objects objects animals animals-objects objects Avg.
SDXL 0.6951 0.8654 0.4926 0.7820 1.5474 0.6789 0.8435
+ resampling 0.7351 0.8666 0.4528 1.0881 1.6755 0.8263 0.9407
   + Corrector 0.7177 0.8794 0.4796 1.0630 1.6429 0.8949 0.9463
   + weight-modulation 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158 1.0118
Table 2: Ablations study conducted on the CO3 method. Starting with the base-model SDXL we progressively add different components.
Comparison with Optimization-Free Approaches: To further demonstrate CO3’s model-agnostic nature and generalization to other base models, we evaluate it on PixArt-$\Sigma$ (Chen et al., 2024). We compare against both the base model and the sampling-based, optimization-free Composable Diffusion (Liu et al., 2022). Unlike Composable Diffusion, which composes diffusion chains within the DDIM prediction step, CO3 applies composition as a corrector. As shown in Fig. 5, Composable Diffusion suffers from missing or mixing concepts, while CO3 consistently preserves all concepts with correct bindings in the output.
Figure 5: Model Agnostic behavior: Qualitative comparison of generation from PixART-$\Sigma$(Chen et al., 2024) base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion.
5 Related Work
Optimization-based correction approaches: (1) Text-embedding optimization works like T2I-Zero variants (Tunanyan et al., 2023), and style or prompt inversion (Li et al., 2023a), improve semantic alignment by optimizing or restructuring the text representation during inference. These methods tweak or invert token embeddings so that the denoising trajectory better reflects entity–attribute bindings or multi-concept prompts.
(2) Attention-map optimization works (Zhang et al., 2024; Meral et al., 2024; Tumanyan et al., 2023; Hertz et al., 2022) directly manipulate the cross- and self-attention maps during sampling. These methods typically inject constraints or losses that (i) boost token–region correspondence for each entity, (ii) reduce overlap between different concepts, or (iii) preserve early layout information across timesteps. Attend-and-Excite (Chefer et al., 2023) increases the activation of object tokens. Divide-and-Bind (Liu et al., 2023) maximizes total variation to elicit distinct excitations for multiple objects and aligns attribute–entity attention maps, while A-STAR (Agarwal et al., 2023) further reduces cross-token overlap and preserves early attention signals. SynGen (Rassin et al., 2023) leverages syntactic parses to penalize mismatched attention overlaps, ensuring linguistic binding between entities and modifiers. InitNO (Guo et al., 2024) optimizes the initial noise so that sampling begins in more favorable regions that yield stronger, less-conflicted attention.
Composable generation works: view conditional diffusion models as energy/score functions, enabling algebraic composition. Composed-Diffusion (Liu et al., 2022) frames score-composition under CFG and demonstrates test-time generalization, but suffers from concept mixing and missing. Follow-ups propose training-free and model-agnostic methods like energy-parameterized diffusion and Metropolis/MCMC-corrected samplers that markedly improve multi-condition generation (Du et al., 2024), but their performance is often poor (Chefer et al., 2023; Feng et al., 2022). Kwon & Ye (2025) uses a composition similar to our resampler but not in the context of a corrector. Instead they use multiple DDIM forward-backward steps to sample initial noise.
Layout augmentation image generation: (1) Layout-to-image methods: (Xie et al., 2023; Phung et al., 2024; Kim et al., 2023; Zhao et al., 2023) use a strategy with explicit spatial priors—-boxes, masks, or region texts to bridge text and image: training-free controllers steer attention so objects materialize in designated regions. Other works add instance-level handles, for fine-grained placement and attributes across multiple entities (Wang et al., 2024). In parallel, fine-tuning approaches inject layout channels into the backbone (Li et al., 2023b; Mou et al., 2023; Zhang et al., 2023).
(2) LLM-augmented methods: Collectively, these works leverage LLM reasoning/representations to better bridge linguistic structure and the denoising trajectory by (i) decomposing complex prompts into regional sub-tasks (Yang et al., 2024; Hu et al., 2024) (ii) infer layouts from text (Qu et al., 2023), and (iii) act as stronger text encoders or timestep-aware adapters  (Saharia et al., 2022b).
6 Conclusion & Future Work
We propose a gradient-free, model-agnostic image composition method that switches between noise-resampling and correction. Our Concept Contrasting corrector is based on the hypothesis that composition is degraded by “problematic” modes—those that overlap with individual concept modes, resulting in strong alignment with particular concepts while suppressing others. Our corrector employs a Closeness-Aware Weight Modulation scheme that emphasizes “pure” modes where all concepts coexist without any dominating ones. Crucially, we attribute shortcomings of past composition methods to the choice of composition weights and show that only when the weights sum to $1$, the interpretation of Tweedie-mean correction and the CFG guidance strength are preserved. Results demonstrate that CO3 outperforms baseline approaches for several diffusion model choices and serves as a plug-and-play module offering improved compositional generation performance.
While our corrector’s closeness-aware weight modulation suppresses the concept mode overlap, the noise/score landscape is strongly biased by the training paradigms of diffusion models, i.e., the quantity and quality of the multi-concept bindings in the training set. This problem also manifests when presented with more unrealistic prompts, resulting in poor text/concept alignment. As shown in Fig. 7, while our work offers a significant boost to multi-concept compositional generation, clearly, there are failure cases that are not yet addressed. However, we believe that more advanced energy-based compositional samplers—that explicitly take into account the probabilities of the landscape during sampling—can further help in avoiding such cases. We hope to investigate these special cases in the future. We also hope to explore applications beyond composition using the CO3 corrector.
7 Reproducibility Statement
All method implementations, inference pipelines are included in the supplementary material. This includes instructions for environment setup, dependencies, and reproducible random seeds. The datasets used in our experiments are publicly available. We provide detailed descriptions of the dataset in Sec. 4. Hyperparameters, inference schedules, and evaluation pipelines are described in Sec. 4, with further details in the Appendix A.3.
Appendix A Appendix
A.1 Broader Impacts
CO3 improves compositional text-to-image generation by adjusting the sampling distribution of diffusion models. At the same time, it introduces certain risks. The method could be misused to create deceptive or misleading visuals, contributing to the spread of misinformation. When applied to depictions of public figures, it may compromise personal privacy. Moreover, the automatically produced content can raise concerns around copyright and intellectual property.
A.2 Composition in Tweedie-denoised space for a valid cfg
Classifier free guidance: In CFG, to sample from $p(x|c)$, we compose the conditonal and unconditonal predicted noise at each time step $t$ as:
${\epsilon}^{\lambda,c}_{t}={\epsilon}^{\phi}_{t}+\lambda({\epsilon}^{c}_{t}-{\epsilon}^{\phi}_{t})$ (16)
where ${\epsilon}^{c}_{t}$ is the noise predicted for the distribution $p_{t}(x_{t}|c)$ at time $t$, i.e., $\nabla_{x_{t}}\log p_{t}(x_{t}|c)=-\frac{{\epsilon}^{c}_{t}}{\sqrt{1-\bar{\alpha}_{t}}}$. Then we use this composed noise to sample the next step as,
$\hat{x}^{\lambda,c}_{t,0}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}{\epsilon}^{\lambda,c}_{t}}{\bar{\alpha}_{t}},\quad x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}^{\lambda,c}_{t,0}+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon^{\lambda,c}_{t}$ (17)
Let, $\hat{x}^{\lambda,c}_{tweedie}:=x_{t}-\sqrt{1-\bar{\alpha}_{t}}{\epsilon}^{\lambda,c}_{t}$ be the tweedie mean from CFG composed noise at time $t$. Then we can rewrite the above equation as,
$x_{t-1}=\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\hat{x}^{\lambda,c}_{tweedie}+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon^{\lambda,c}_{t}$ (18)
Composition of CFG scores: (Liu et al., 2022) introduced the idea composable-diffusion where they compose scores from different diffusion models or conditional distributions to generate images from the composed distribution. They assume that, $p(x_{0}|C)=\prod p(x_{0}|c_{i})$ and proposed the modified CFG composition for each $t$ as,
$\tilde{{\epsilon}}^{\lambda,C}_{t}={\epsilon}^{\phi}_{t}+\lambda_{i}(\sum_{i}{\epsilon}^{c_{i}}_{t}-{\epsilon}^{\phi}_{t})$ (19)
The issue with above composition is that for arbitrary $\lambda_{i}$, the composed CFG noise doesn’t satisfy the CFG equation in equation 16, i.e., $\tilde{{\epsilon}}^{\lambda,C}_{t}\neq{\epsilon}^{\lambda,C}_{t}$ for any $\lambda$ in equation equation 16 where ${\epsilon}^{\lambda,C}_{t}$ is the noise predicted for the distribution $p_{t}(x_{t}|C)$ at time $t$.
To conretize this, we state the following simple result.
Lemma 2.
Let’s define CFG as a function at time $t$ as, $f_{CFG}({\epsilon};\lambda)=$ ${\epsilon}^{\phi}+\lambda({\epsilon}-{\epsilon}^{\phi})$. Then for any $\lambda$ and $K>1$, $\sum_{k}{w_{k}f_{CFG}}=f_{CFG}(\sum_{k}{w_{k}{\epsilon}^{k}})$ only if $\sum_{k}w_{k}=1$.
Proof.
Let, $\tilde{{\epsilon}}^{\lambda}_{t}=\sum_{k}{w_{k}f_{CFG}}$. Then,
$\tilde{{\epsilon}}^{\lambda}_{t}$ $=\sum_{k}w_{k}({\epsilon}^{\phi}+\lambda({\epsilon}^{k}-{\epsilon}^{\phi}))$
$=\sum_{k}w_{k}{\epsilon}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-\sum_{k}w_{k}{\epsilon}^{\phi})$
$={\epsilon}_{\phi}\sum_{k}w_{k}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-{\epsilon}^{\phi}\sum_{k}w_{k})$
$={\epsilon}_{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-{\epsilon}^{\phi})\quad\text{if }\sum_{k}w_{k}=1$
$=f_{CFG}(\sum_{k}w_{k}{\epsilon}^{k})\quad\text{if }\sum_{k}w_{k}=1$
If $\sum_{k}w_{k}\neq 1$, then $\tilde{{\epsilon}}^{\lambda}_{t}\neq f_{CFG}(\sum_{k}w_{k}{\epsilon}^{k})$. ∎
Remark: One immediate implication of this lemma is that, whenever we sample from a composed distribution with the CFG composed score of the form $\tilde{{\epsilon}}^{\lambda}=w_{1}{\epsilon}^{\lambda,1}+w_{2}{\epsilon}^{\lambda,2}+\dots+w_{K}{\epsilon}^{\lambda,K}$, then $\tilde{{\epsilon}}^{\lambda}$ corresponds to the CFG noise for the distribution $p(x|C)=\prod_{i}p(x\mid c_{i})^{w_{i}}$ only when $\sum_{i}w_{i}=1$.
Lemma 3.
Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$. Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$. Then,
  1. a)
    CO3-corrector: $\tilde{\hat{x}}_{tweedie}$ can be expressed in the form of a tweedie-mean at time $t$, i.e. $\tilde{\hat{x}}_{tweedie}=x-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ if and only if $\sum_{k}w_{k}=1$. Here $\tilde{\lambda}=\lambda$ and CFG composed noise $\tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$.
  2. b)
    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
Proof.
a) The proof can be obtained by simplifying the $\hat{x}^{comp}_{tweedie}$ expression and using direct application of Lemma 2. We have,
$\tilde{\hat{x}}_{tweedie}$ $=\sum_{k}w_{k}\hat{x}^{\lambda,c_{k}}_{tweedie}$
$=\sum_{k}w_{k}(x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c_{k}})$
$=x_{t}\sum_{k}w_{k}-\sigma_{t}\sum_{k}w_{k}{\epsilon}_{t}^{\lambda,c_{k}}$
[…]
$=x_{t}-\sigma_{t}\ f_{CFG}(\sum_{k}w_{k}{\epsilon}^{c_{k}})\quad\text{if }\sum_{k}w_{k}=1$
So, we have $\tilde{\hat{x}}_{tweedie}=x_{t}-\sigma_{t}\ \tilde{{\epsilon}_{t}}^{\lambda,C}$ where $\tilde{{\epsilon}_{t}}^{\lambda,C}=f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}})={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$. If $\sum_{k}w_{k}\neq 1$, then $\tilde{\hat{x}}_{tweedie}$ cannot be expressed in the form $x_{t}-\sigma_{t}\ \tilde{{\epsilon}_{t}}^{\lambda,C}$ for any $\lambda$ and $\tilde{{\epsilon}_{t}}^{\lambda,C}$.
b) From the $4$-th equality of part a)
$\tilde{\hat{x}}_{tweedie}$ $=x_{t}\sum_{k}w_{k}-\sigma_{t}\sum_{k}w_{k}{\epsilon}_{t}^{\lambda,c_{k}}$
[…]
A.3 More Implementation Details
A.3.1 CO3 Implementation Details
For the results in Table 1 we use $T_{c}=10$(number of time-steps to correct), $T_{r}=3$(number of resampling steps) and P=5P=5(number of corrector iteration). For concept aware weight modulation, we use exponential kernel with $\beta=0.8$ while anchoring $w_{0}$ at $1.0$ for CO3-resampler and $2.0$ for CO3-corrector.
We use Stanza (Qi et al., 2020) to parse the prompts. We parse the prompts to extract different noun chunks and filter each of them to remove articles and adjectives. The remaining proper noun is used as concept in CO3. For example, if $C$ is ”a black cat and a brown dog”, we consider $c_{1}=$”cat” and $c_{2}=$”dog”.
A.3.2 Baseline Methods
Attend-Excite (Chefer et al., 2023), Divide-Bind (Liu et al., 2023), InitNo (Guo et al., 2024), SynGen (Rassin et al., 2023) and Composable Diffusion (Liu et al., 2022) are methods based on SD1.5. We run experiments on these models using their publicly available code. ToMe (taihang Hu et al., 2024) is SDXL based we use their publicly available code-base implementation. We also adapted Composable Diffusion and SynGen to SDXL.
A.4 Time Complexity Analysis
We conduct an experiment to analyze the time-complexity of each SDXL based methods. The Table 3 indicates the average time (in secs) taken by the model per sample. The experiments are conducted on NVIDIA A100 gpu.
Method Inference Steps Time Cost (sec) Animals Animals-Objects Objects
SDXL 50 7.20 0.6950 0.8654 0.4926
Composable Diffusion 50 9.80 0.2846 0.5656 0.4529
SynGen(SDXL) 50 11.48 0.6816 0.8578 0.4652
ToMe 50 16.58 0.6257 0.8808 0.6440
CO3(ours) 50 19.9 0.7441 0.8878 0.5146
Table 3: Comparison of different methods across efficiency and compositionality metrics.
A.5 Comparison of compositions in score space
In this section we compare our composition framework with Composable-Diffusion. As already described in section 3, we propose composition in Tweedie-denoised space as
$\tilde{x}_{tweedie}=w_{0}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,C}]\;+\;w_{1}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{1}}]\;+\;\dots\;+\;w_{K}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{K}}]$ (20)
which leads to
$\tilde{x}_{tweedie}$ $=x-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}\text{ iff }\sum w_{k=0}^{K}=1$ (22)
$\text{ where }\tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ $={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$ (23)
Contrast this with the Composable-Diffusion’s noise/score composition:
$\tilde{\epsilon}^{\lambda,C}_{t,compdiff}\;=\;\epsilon_{t}^{\phi}+\lambda_{1}\bigl(\epsilon_{t}^{c_{1}}-\epsilon_{t}^{\phi}\bigr)+\lambda_{2}\bigl(\epsilon_{t}^{c_{2}}-\epsilon_{t}^{\phi}\bigr)+\dots+\lambda_{K}\bigl(\epsilon_{t}^{c_{K}}-\epsilon_{t}^{\phi}\bigr)$ (24)
For arbitrary weights $\lambda_{k}$}, this cannot be expressed in the from equation 23. This proves that $\tilde{\epsilon}^{\lambda,C}_{t,compdiff}$ doesn’t lead to a valid Tweedie-mean.
A.6 More Visualizations
Figure 6: Qualitative results of SDXL base diffusion model, and SDXL + CO3 for animal-animal,animal-object, and object-object prompt categories
A.6.1 CO3 Failure Cases
Despite correcting for the “problematic” modes, there still remain open challenges in image composition as shown in Figure 7. The score and, hence, the correction landscape is heavily influenced by the training schemes employed in diffusion model training, i.e., the quantity and quality of the multi-concept bindings in the training set. In addition, the usage of unrealistic prompts perhaps not encountered in training also results in poor text/concept alignment. We leave this investigation for the future.
Figure 7: Failure scenarios of PixART-$\Sigma$ base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion.
A.7 Ablations on Hyperparameters
In this section, we analyze the contribution of the following five factors to the performance of our CO3 corrector.
Notation recap: $\beta$ is the exponential decay factor in the affinity scores; $\lambda$ scales the composed_score for CFG; num_resampling is the number of resampling steps at the start of diffusion; num_ts is the number of early timesteps where the corrector is used; num_steps is the number of iterations per corrector application.
A.7.1 Number of resampling steps
Table 4 shows results for different resampling steps. We observe that using more resampling steps can blur concept separation early in the diffusion process rather than helping to disentangle concepts.
Table 4: num_resampling sweep. Within each group, only num_resampling changes; all other settings are identical.
Frozen settings num_resampling ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects $\beta=0.8$, num_ts=7, $\lambda=0.8$, num_steps=5 3 1.2218 1.7020 0.9405 0.7375 0.8811 0.4782 4 1.1705 1.7031 0.9512 0.7259 0.8828 0.4696 $\beta=0.3$, num_ts=7, $\lambda=0.8$, num_steps=5 3 1.1484 1.6915 0.9005 0.7240 0.8810 0.4691 4 1.1297 1.6948 0.8996 0.7145 0.8774 0.4697 $\beta=1.1$, num_ts=7, $\lambda=0.8$, num_steps=5 3 1.2566 1.7062 0.9916 0.7422 0.8813 0.4834 4 1.1927 1.7164 0.9934 0.7292 0.8835 0.4811
A.7.2 Corrector application timesteps
In CO3, the corrector is applied during the first num_ts diffusion steps. Table 5 shows that larger num_ts generally yields better results, especially for animals and objects. This indicates the corrector remains beneficial beyond only the earliest timesteps.
Table 5: num_ts sweep with all other settings fixed: $\beta=1.1$, num_resampling=3, num_steps=5, $\lambda=0.8$. (Best values are bold)
num_ts ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects 4 1.1862 1.7136 0.8885 0.7430 0.8783 0.4630 6 1.2510 1.7117 0.9750 0.7432 0.8801 0.4812 7 1.2566 1.7062 0.9916 0.7422 0.8813 0.4834 10 1.3149 1.6816 1.0095 0.7576 0.8832 0.5074
A.7.3 Corrector iterations
Table 6 reports the effect of increasing the number of corrector iterations. Despite setting num_ts=4, raising num_steps reduces performance substantially, suggesting that applying the corrector only at the very beginning is not sufficient; employing it until much later in the diffusion trajectory is more effective.
Table 6: num_steps. Within each group, only num_steps changes; all other settings are identical.
Fixed settings num_steps ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects (a) $\beta=1.1$, num_ts=4, num_resampling=2, $\lambda=0.9$ 7 1.2622 1.7129 0.9959 0.7542 0.8843 0.4707 10 1.2452 1.6994 0.9753 0.7448 0.8809 0.4706 15 1.2151 1.6701 0.9437 0.7318 0.8781 0.4556 (b) $\beta=1.1$, num_ts=4, num_resampling=2, $\lambda=0.8$ 7 1.2755 1.7049 0.9424 0.7538 0.8762 0.4611 10 1.2503 1.6969 0.9191 0.7446 0.8798 0.4617 15 1.2480 1.6658 0.9063 0.7356 0.8750 0.4577
A.7.4 Exponential decay factor $\beta$ of affinity scores
We vary $\beta$ in  equation 14 from 0.3 to 1.1. Tables 7 and 8 demonstrate that the performance generally increases with larger $\beta$, with the strongest gains in the animals and objects categories, while a few settings exhibit minor regressions.
Table 7: Exponential decay factor $\beta$ with all other settings fixed: num_resampling=3, num_steps=5, $\lambda=0.8$. We report two blocks: num_ts=6 and num_ts=7. (Best values for each block are bold)
num_ts $\beta$ ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects 6 0.3 1.1458 1.7016 0.8801 0.7262 0.8793 0.4613 0.6 1.1945 1.7069 0.9010 0.7317 0.8786 0.4645 0.7 1.2043 1.7109 0.9152 0.7310 0.8800 0.4706 0.8 1.2157 1.7024 0.9156 0.7398 0.8776 0.4751 0.9 1.2305 1.7146 0.9511 0.7400 0.8809 0.4766 1.0 1.2282 1.7183 0.9596 0.7350 0.8808 0.4770 1.1 1.2510 1.7117 0.9750 0.7432 0.8801 0.4811 7 0.3 1.1484 1.6915 0.9005 0.7240 0.8809 0.4691 0.8 1.2218 1.7020 0.9405 0.7375 0.8811 0.4782 1.1 1.2566 1.7062 0.9916 0.7422 0.8813 0.4834
Table 8: Exponential decay factor $\beta$ with all other settings fixed: num_resampling=3, num_steps=5, $\lambda=0.9$. We report two blocks: num_ts=6 and num_ts=7. (Best values for each block are bold.)
num_ts $\beta$ ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects 6 0.3 1.1437 1.7020 0.9518 0.7248 0.8816 0.4764 0.7 1.2051 1.7130 0.9860 0.7407 0.8851 0.4788 0.8 1.2037 1.7110 0.9916 0.7388 0.8832 0.4894 0.9 1.2003 1.7159 0.9883 0.7416 0.8842 0.4891 1.1 1.2098 1.7180 0.9956 0.7371 0.8841 0.4855 7 0.3 1.1273 1.6938 0.9490 0.7274 0.8848 0.4796 0.8 1.2142 1.7000 0.9937 0.7363 0.8845 0.4952 1.1 1.2331 1.7046 1.0148 0.7496 0.8864 0.4945

### v2

@@
-CO3: Contrasting Concepts Compose Better
+Steer Away from Mode Collisions: Improving Composition in Diffusion Models
 Abstract
@@
 Instead of re-training, we introduce a corrective sampling strategy that steers away from regions where the joint prompt behavior overlaps too strongly with any single concept in the prompt.
-The goal is to steer towards “pure” joint modes where all concepts can coexist with balanced visual presence..
+The goal is to steer towards “pure” joint modes where all concepts can coexist with balanced visual presence.
 We further show that existing multi‑concept guidance schemes can operate in unstable weight regimes that amplify imbalance; we characterize favorable regions and adapt sampling to remain within them.
@@
 Results suggest that lightweight corrective guidance can substantially mitigate brittle semantic alignment behavior in modern diffusion systems.
+Code is available at https://github.com/debottam-dutta7/co3
 1 Introduction
@@
 However, ensuring the generated images faithfully adhere to the prompt, a challenge known as semantic alignment (Chefer et al., 2023; taihang Hu et al., 2024; Liu et al., 2023), remains a problem.
-Concretely, for a given prompt $C$, T2I models like StableDiffusion (Rombach et al., 2022) sample from the modes (or high probability regions) of the learned distribution, $p(x|C)$.
+Concretely, for a given prompt $C$, T2I models like StableDiffusion (Rombach et al., 2022) sample from the modes (or high probability regions) of the learned distribution, $p(x\mid C)$.
 While such models can produce high resolution images in general, every so often, the results are surprisingly misaligned even for very simple prompts containing few concepts, e.g., CC=“a cat and a dog”.
 Diagnosing exactly why this behavior emerges periodically is difficult.
-It is conceivable that the complex training process in high dimensions, especially in conjunction with text embeddings, creates some problematic modes in $p(x|C)$.
-We hypothesize that problematic modes in $p(x|C)$ arise when they overlap with modes of individual concept distribution $p(x|c_{i})$.
+It is conceivable that the complex training process in high dimensions, especially in conjunction with text embeddings, creates some problematic modes in $p(x\mid C)$.
+We hypothesize that problematic modes in $p(x\mid C)$ arise when they overlap with modes of individual concept distribution $p(x\mid c_{i})$.
 Such an overlap biases the generation toward a single concept $c_{i}$, reducing the prominence of others.
 For instance, across images of $c_{1}=\text{``cat''}$ in the training dataset, a few may have an inconspicuous or partial $c_{2}=\text{``dog''}$ in the background.
-This image may still fall under the mode of $p(x|C)$.
+This image may still fall under the mode of $p(x\mid C)$.
 We attribute this to training instabilities and relatively less coverage of multi-concept prompts $C$, which cause the model to assign high probability even to weakly conforming images.
-Said differently, an image of a big cat and an inconspicuous dog can get assigned high probabilities under $p(x|C)$, causing semantic misalignment.
+Said differently, an image of a big cat and an inconspicuous dog can get assigned high probabilities under $p(x\mid C)$, causing semantic misalignment.
 Preventing such problematic modes warrants strict and specialized training paradigms; a difficult task for such large models.
@@
 Figure 1 illustrates the intuition behind our proposal.
-Our corrector distribution $\tilde{p}(x|C)$ assigns low probability to regions where $p(x|C)$ overlaps with individual $p(x|c_{i})$; we deem them as degenerate modes dominated by a single concept.
-By suppressing these overlaps, the corrector emphasizes pure $p(x|C)$ modes where all concepts coexist without one overwhelming the others.
-From a probabilistic perspective, this acts as a corrective factor: while $p(x|C)$ may assign high probability to weakly conforming images due to training noise or limited multi-concept data, dividing by the marginals removes this bias and sharpens the distribution toward genuine multi-concept samples.
+Our corrector distribution $\tilde{p}(x\mid C)$ assigns low probability to regions where $p(x\mid C)$ overlaps with individual $p(x\mid c_{i})$; we deem them as degenerate modes dominated by a single concept.
+By suppressing these overlaps, the corrector emphasizes pure $p(x\mid C)$ modes where all concepts coexist without one overwhelming the others.
+From a probabilistic perspective, this acts as a corrective factor: while $p(x\mid C)$ may assign high probability to weakly conforming images due to training noise or limited multi-concept data, dividing by the marginals removes this bias and sharpens the distribution toward genuine multi-concept samples.
 As a result, the modes we target are more semantically aligned and less prone to concept suppression or distortion.
-Correction sampling from $\tilde{p}(x|C)$ can be achieved by composing scores from constituent component distributions $\nabla_{x}\log p(x|C),\{\nabla_{x}\log p(x|c_{i})\}_{i}$ (Liu et al., 2022).
+Correction sampling from $\tilde{p}(x\mid C)$ can be achieved by composing scores from constituent component distributions $\nabla_{x}\log p(x\mid C),\{\nabla_{x}\log p(x\mid c_{i})\}_{i}$ (Liu et al., 2022).
 While there are many ways to compose, we analyse and show that composition through weighted sum of Tweedie-means—in the Tweedie denoised space—offers a more general framework that subsumes existing approaches.
-In particular, we show that two classes of correction sampling—noise-resampling and latent correction (Bao et al., 2024; Rassin et al., 2023; taihang Hu et al., 2024; Kwon & Ye, 2025)—become special cases of Tweedie-mean composition under different weighting schemes.
+In particular, we show that two classes of correction sampling—noise-resampling and latent correction (Bao et al., 2024; Rassin et al., 2023; taihang Hu et al., 2024; Kwon and Ye, 2025)—become special cases of Tweedie-mean composition under different weighting schemes.
 This allows us to design a hybrid composition framework that serves the purpose of resampling at time $T$, and then toggles to latent correction at later steps.
-By latent, we mean that the correction steps are accomplished in between each DDIM time step, allowing CO3 to be a plug-and-play, model-agnostic and gradient-free approach for T2I models.
+By latent, we mean that the correction steps are accomplished in between each DDIM time step, allowing CO3(Concept Contrasting Corrector) to be a plug-and-play, model-agnostic and gradient-free approach for T2I models.
 Comparison of CO3 against SOTA baselines shows stronger semantic alignment to prompts (measured using multiple metrics), giving empirical evidence of our hypothesis.
 Figure 1: The figure illustrates our hypothesis on mode overlap using a simple 2D toy example.
-(a) Two modes of the distribution $p_{t}(x|\textit{"a cat and a dog"})$ (in green contour) has significant overlap with the modes of the individual concept distributions $p_{t}(x|\textit{"a cat"})$ (in red contour) and $p_{t}(x|\textit{"a dog"})$ (in orange contour).
-(b) The proposed corrector distribution $p_{t}(x|\texttt{"a cat and a dog"})/(p_{t}(x|\textit{"a cat"})p_{t}(x|\textit{"a dog"}))$ suppresses these overlaps, steering the generation away from problematic modes.
+(a) Two modes of the distribution $p(x\mid\texttt{"a cat and a dog"})$ (in green contour) has significant overlap with the modes of the individual concept distributions $p(x\mid\texttt{"a cat"})$ (in red contour) and $p_{t}(x\mid\texttt{"a dog"})$ (in orange contour).
+(b) The proposed (unnormalized) corrector distribution $p(x\mid\texttt{"a cat and a dog"})/(p(x\mid\texttt{"a cat"})p(x\mid\texttt{"a dog"}))$ suppresses these overlaps, steering the generation away from problematic modes.
 The arrows indicate the denoising directions.
@@
 This perspective highlights how the Tweedie decomposition denoising via $\hat{x}_{0}$, followed by deterministic reconstruction of $x_{t-1}$, is naturally aligned with DDIM sampling.
-In practice, most T2I models adopt classifier-free guidance (CFG) (Ho & Salimans, 2022), where predictions from both the conditional and unconditional models are combined:
+In practice, most T2I models adopt classifier-free guidance (CFG) (Ho and Salimans, 2022), where predictions from both the conditional and unconditional models are combined:
 $\epsilon_{t}^{\lambda,c}\;=\;\lambda\ \epsilon_{\theta}(x_{t},c,t)\;+\;(1-\lambda)\,\epsilon_{\theta}(x_{t},\varnothing,t)$ (4)
@@
 where $\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}]:=x_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}^{\lambda,c}$ is the Tweedie mean from the CFG noise at $t$.
-Thus, CFG in the Tweedie framework can be interpreted as modifying the denoised estimate $\hat{x}_{0}$ to $\hat{x}_{0}^{\lambda,c}$ before the deterministic step to $x_{t-1}$ (Chung et al., 2024; Kwon & Ye, 2023).
+Thus, CFG in the Tweedie framework can be interpreted as modifying the denoised estimate $\hat{x}_{0}$ to $\hat{x}_{0}^{\lambda,c}$ before the deterministic step to $x_{t-1}$ (Chung et al., 2024; Kwon and Ye, 2023).
 $\blacksquare$ Correction-based approaches for conditional generation.
@@
 $\blacksquare$ Composable-diffusion. Generating samples that satisfy multiple conditions $\{c_{1},\ldots,c_{K}\}$ can be formulated as sampling from the joint distribution
-$\tilde{p}_{0}(x_{0}\mid c_{1},\ldots,c_{K})\;\propto\;\prod_{k=1}^{K}p_{0}(x_{0}\mid c_{k}).$ (9)
+$\tilde{p}_{0}(x_{0}\mid c_{1},\ldots,c_{K})\;\propto\;p(x_{0})\prod_{k=1}^{K}p_{0}(c_{k}\mid x_{0}).$ (9)
 To achieve this, Liu et al. (2022) proposed composable diffusion, which directly composes the output scores (predicted noises) from different conditional diffusion models using CFG during sampling.
@@
 In summary, both the correction-based approach and the composable diffusion idea can be interpreted as different ways of approximating Tweedie-means $\hat{x}_{tweedie}[\epsilon^{\lambda,c}_{t}]$ at time $t$.
-3 CO3: Contrasting Concepts to Compose
+3 CO3: Concept Contrasting Corrector
 We aim to combine the strengths of correction-based approaches and composable diffusion.
 On one hand, correction-based methods are powerful and explicitly improve compositional alignment, but they are subject to the complex gradient of the base model.
-On the other hand, composable diffusion is fully model-agnostic, but suffers from poor performance because linear score composition is not consistent with the diffusion forward process.
+On the other hand, composable diffusion is fully model-agnostic, but suffers from poor performance because linear score composition is not consistent with the diffusion forward process Du et al.
+(2024).
 To take advantage of both, assume that prompt $C$ can be decomposed into constituent concepts $\{c_{1},c_{2},\ldots,c_{K}\}$.
-We propose an explicit Concept Contrasting corrector distribution based on our hypothesis on mode overlap discussed in Sec.
+We propose an explicit Concept Contrasting Corrector (CO3) distribution based on our hypothesis on mode overlap discussed in Sec.
 1.
@@
 7:  end if
-8:  $x_{t}\leftarrow\text{DDIM\_step}(x_{t})$ $\triangleright$ Inverse DDIM
+8:  $x_{t}\leftarrow\text{DDIM\_step}(x_{t})$ $\triangleright$ Reverse DDIM
 9:end for
@@
 13 must be one that can be expressed in the definition given in equation 5 as DDIM goes through a series of Tweedie means for image generation as described in section 2.
-Lemma 1 shows that this is guaranteed only when the weights satisfy the normalization condition $\sum_{i}w_{i}=1$.
-Lemma 1.
-Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$.
+Proposition 1 shows that this is guaranteed only when the weights satisfy the normalization condition $\sum_{i}w_{i}=1$.
+Proposition 1.
+Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise ${\epsilon}_{t}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$.
 Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$.
@@
     CO3-corrector: $\tilde{\hat{x}}_{tweedie}$ can be expressed in the form of a tweedie-mean at time $t$, i.e.
-$\tilde{\hat{x}}_{tweedie}=x-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ if and only if $\sum_{k}w_{k}=1$.
+$\tilde{\hat{x}}_{tweedie}=x_{t}-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ if and only if $\sum_{k}w_{k}=1$.
 Here $\tilde{\lambda}=\lambda$ and CFG composed noise $\tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$.
   2. b)
-    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
+    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
 […]
@@
 so that concepts closer to the current sample (smaller $d_{k}$) receive higher affinity. These affinity scores are then normalized to define the weights:
-$w_{k}=-\frac{a_{k}}{\sum_{j=1}^{K}a_{j}},$ (15)
+$\vskip-2.84544ptw_{k}=-\frac{a_{k}}{\sum_{j=1}^{K}a_{j}},\vskip-2.84544pt$ (15)
 which ensures each weight is negative and the total sum satisfies $\sum_{k=1}^{K}w_{k}=-1$.
@@
 All experiments are performed on a single NVIDIA A6000 GPU (48GB).
-Additional hyperparameter details are provided in the Appendix section A.3.
-Evaluation benchmark and metrics: We evaluate our method on the benchmark dataset from Chefer et al.
-(2023), which includes three categories of prompts: Animal–Animal, Animal–Object, and Object–Object.
-While the benchmark is limited to these cases, CO3 naturally extends to more complex prompts involving arbitrary numbers of concepts.
-For evaluation, we adopt BLIP-VQA (Huang et al., 2023) and ImageReward (Xu et al., 2023), both widely used in recent works (taihang Hu et al., 2024; Hu et al., 2024) to measure the faithfulness of generated images to input prompts.
-Notably, ImageReward is a learned reward model trained on human preferences, capturing aspects such as prompt adherence, aesthetics, and overall quality.
-For both metrics, higher scores indicate better alignment.
-Comparison methods: We compare CO3 against state-of-the-art approaches, including optimization-based correction methods (Attend-and-Excite (Chefer et al., 2023), SynGen (Rassin et al., 2023), Divide-and-Bind (Liu et al., 2023), ToME (taihang Hu et al., 2024)), the composable generation method Composable Diffusion (Liu et al., 2022), and the noise-optimization method InitNO (Guo et al., 2024).
+Additional implementation details are provided in the Appendix section A.4.
+Evaluation benchmark and metrics: We first evaluate our method on simpler prompts (two concepts) using the Attend-Excite benchmark from Chefer et al.
+(2023) which contains categories of prompts: Animal–Animal, Animal–Object, and Object–Object.
+We then take the best performing baselines from this task and compare them with CO3 on more complex prompts (three or more concepts) from T2ICompbench Huang et al.
+(2023) and rare or novel concept prompts from RareBench (Park et al., 2024).
+Specifically, we consider ”Complex” category from T2ICompbench and ”Concat”, ”Relations” and ”Complex” from RareBench; the subset of categories which stress multi-concept interactions, attribute entanglement and rare-concept scenarios.
+For evaluation, we adopt BLIP-VQA (Huang et al., 2023), ImageReward (Xu et al., 2023) and DINOv2 (Oquab et al., 2023) scores for simpler prompts.
+While BLIP-VQA and ImageReward (widely used in recent works (taihang Hu et al., 2024; Hu et al., 2024)) measure the faithfulness of generated-images to input-prompts, DINOv2 measures the quality of concepts generated by each model via a classification score.
+R2F Park et al.
+(2024) notes that BLIP-VQA can be unreliable in rare-concept scenarios; hence, following their recommendation, we report metrics that more closely reflect human judgement, namely, ImageReward and human evaluation (from 11 participants).
+Comparison methods: We compare CO3 against state-of-the-art approaches, including optimization-based correction methods (Attend-Excite (Chefer et al., 2023), SynGen (Rassin et al., 2023), Divide-Bind (Liu et al., 2023), Magnet Zhuang et al.
+(2024), ToME (taihang Hu et al., 2024)); composable generation methods Comp-Diff (Liu et al., 2022), Tweediemix (Kwon and Ye, 2025); LLM augmented method R2F (Park et al., 2024) and the noise-optimization method InitNO (Guo et al., 2024).
+For fair comparison against Tweediemix’s sampling strategy, we implemented Tweediemix without its LORA-finetuned customization stage.
 All optimization-based approaches are architecture-dependent and require gradient computations at inference, with InitNO being model-agnostic but still incurring costly backpropagation.
-In contrast, only Composable Diffusion and our proposed CO3 are purely sampling-based, model-agnostic, and completely gradient-free.
+In contrast, only Composable Diffusion, Tweediemix and CO3 are purely sampling-based, model-agnostic, and completely gradient-free.
 4.2 Experimental Results
-Properties BLIP-VQA ImageReward
-training-free gradient-free model-agnostic animals animals-objects objects animals animals-objects objects
-SD1.5(Rombach et al., 2022) - 0.3239 0.5958 0.2730 -0.2733 0.4262 -0.5521
-Attend-excite(Chefer et al., 2023) 0.6980 0.7865 0.5155 0.8244 1.2380 0.8741
-SynGenRassin et al. (2023) 0.3348 0.7689 0.6595 -0.2445 0.9838 0.7315
-Divide & Bind 0.7201 0.8399 0.5887 0.8499 1.2516 0.8134
-InitNO(Guo et al., 2024) 0.7264 0.7998 0.5406 1.0082 1.3927 1.1383
-SDXLPodell et al. (2023) - 0.6950 0.8654 0.4926 0.7820 1.5574 0.6789
-SynGen(SDXL)Rassin et al. (2023) 0.6816 0.8578 0.4652 0.6998 1.5622 0.6441
-ToMe(taihang Hu et al., 2024) 0.6257 0.8808 0.6440 0.3895 1.5736 1.0118
-Compose-diffusion(Liu et al., 2022) 0.2846 0.5656 0.4529 -1.1399 -0.2068 -0.0955
-CO3(Ours) 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158
-Table 1: Quantitative comparison of different methods on compositional generation tasks.
-We evaluate the generated images using two metrics: BLIP-VQA and ImageReward.
-Top performing model is highlighted in Black and 2nd best in Blue.
-Higher the score the better.
+Table 1: Quantitative comparison of different methods on two concept prompts from Chefer et al.
+(2023).
+We evaluate the generated images using three metrics: BLIP-VQA, ImageReward and DINOv2 scores on categories Animal-Animal(A-A), Animal-Object(A-O) and Object-Object(O-O).
+The top performing model is highlighted in and the 2nd best in .
+Properties BLIP-VQA $\uparrow$ ImageReward $\uparrow$ DINOv2 $\uparrow$
+train-free grad-free model-agnostic A-A A-O O-O A-A A-O O-O A-A A-O O-O
+SD1.5(Rombach et al., 2022) - 0.3239 0.5958 0.2730 -0.2733 0.4262 -0.5521 0.2163 0.2651 0.3010
+Attend-excite(Chefer et al., 2023) 0.6980 0.7865 0.5155 0.8244 1.2380 0.8741 0.2245 0.2599 0.2921
+SynGenRassin et al. (2023) 0.3348 0.7689 0.6595 -0.2445 0.9838 0.7315 0.2187 0.2606 0.2790
+Divide & Bind 0.7201 0.8399 0.5887 0.8499 1.2516 0.8134 0.2029 0.2481 0.2865
+InitNO(Guo et al., 2024) 0.7264 0.7998 0.5406 1.0082 1.3927 1.1383 0.2252 0.2654 0.2953
+SDXLPodell et al. (2023) - 0.6950 0.8654 0.4926 0.7820 1.5574 0.6789 0.2234 0.2618 0.2629
+Magnet Zhuang et al. (2024) 0.6782 0.8744 0.5398 0.6450 1.5895 0.7856 0.2122 0.2636 0.2624
+SynGen(SDXL)Rassin et al. (2023) 0.6816 0.8578 0.4652 0.6998 1.5622 0.6441 0.2221 0.2650 0.2710
+ToMe(taihang Hu et al., 2024) 0.6257 0.8808 0.6440 0.3895 1.5736 1.0118 0.2226 0.2620 0.2545
+TweediemixKwon and Ye (2025)111adapted without the (LORA finetuned) concept-customization stage. 0.7390 0.8059 0.4683 1.0023 1.3127 0.7959 0.2519 0.3105 0.3237
+Comp-diff(Liu et al., 2022) 0.2846 0.5656 0.4529 -1.1399 -0.2068 -0.0955 0.2497 0.2756 0.2908
+CO3(Ours) 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158 0.2623 0.3259 0.3283
+Table 2: Quantitative results on multi-concept prompts from T2ICompbench and rare-concepts from RareBench. We report ImageReward (IR) and Human Evaluation (HE) scores.
+T2ICompbench RareBench
+Complex Concat Relations Complex
+Method IR $\uparrow$ HE $\uparrow$ IR $\uparrow$ HE $\uparrow$ IR $\uparrow$ HE $\uparrow$ IR $\uparrow$ HE $\uparrow$
+SDXL (Podell et al., 2023) 0.3083 0.5983 0.1018 0.5416 0.3607 0.5016 1.1907 0.6283
+SynGen(SDXL) (Rassin et al., 2023) 0.2889 0.5133 0.0576 0.3177 1.1452
+ToMe (taihang Hu et al., 2024) 0.2891 0.5866 -0.2041 0.4500 -0.0632 0.4983 1.0200 0.48666
+R2F (Park et al., 2024) 0.3561 0.4816 -0.0162 0.4300 0.3620 0.5200 1.1542 0.53666
+CO3(Ours) 0.4406 0.8033 0.1610 0.6167 0.2588 0.6200 1.2211 0.61166
 Quantitative Comparison:
-Table 1 demonstrates the extent to which CO3, despite being model-agnostic and optimization-free, competes or outperforms optimization-based and model-agnostic baselines on ImageReward.
+Simpler prompts (Attend-Excite): Table 1 demonstrates the extent to which CO3, despite being model-agnostic and optimization-free, competes or outperforms optimization-based and model-agnostic baselines on ImageReward.
 Gains are substantive in the Animal–Animal and Animal–Object categories.
-On BLIP-VQA, CO3 delivers the best performance in these two categories and improves over the base SDXL on the Object category, which contains prompts of the form: “a [color1] [object1] and a [color2] [object2].”
-Unlike methods such as ToMe taihang Hu et al.
-(2024), SynGen (Rassin et al., 2023), Attend-and-Excite (Chefer et al., 2023), and Divide-and-Bind Liu et al.
+On BLIP-VQA, CO3 delivers the best performance in these two categories and improves over the base SDXL on the Object category, which contains prompts of the form: “a [color1] [object1] and a [color2] [object2].” Unlike methods such as ToMe taihang Hu et al.
+(2024), SynGen (Rassin et al., 2023), Attend-Excite (Chefer et al., 2023), and Divide-Bind Liu et al.
 (2023), which rely on model-specific architectures or explicit attention-based binding losses, CO3 is architecture-independent, sampling-based, and free of explicit subject–attribute binding.
 Interestingly, by simply targeting “pure” modes under $p(x\mid C)$, CO3 discovers layouts that better capture multi-object relations (Animals–Animals, Animal–Objects) and finds high-probability regions that preserve subject–attribute bindings (Objects).
-This performance gain empirically validates our hypothesis on mode overlap and also highlights CO3’s effectiveness in correcting the compositional limitations of SDXL.
-Qualitative Comparison: Figure 3 compares CO3 with several baselines.
-Columns 1–2 show results from Divide-Bind (Liu et al., 2023) and Attend-and-Excite (Chefer et al., 2023) (based on SDv1.5 (Rombach et al., 2022)), while the remaining columns compare against Composable Diffusion (Liu et al., 2022), SynGen (Rassin et al., 2023), and ToMe (taihang Hu et al., 2024).
+Figure 3: Qualitative comparison of different methods on simpler prompts.
+Multiple and rare-concept prompts (T2ICompbench and RareBench): Table 2 shows that CO3 consistently outperforms or matches the best-performing baselines across all categories and metrics.
+We note that CO3 is not designed specifically for rare-concept scenarios—it is a general compositional sampling correction that does not rely on rarity-aware priors or LLM-based prompt augmentation.
+Yet, CO3 matches or exceeds R2F on RareBench, despite R2F being tailored for rare-concept generation.
+We believe this indicates that the early-alignment phase of CO3, which mitigates concept dominance and encourages sampling from “pure” modes, is naturally beneficial for rare-concept generation as well.
+These performance gains empirically validates our hypothesis on mode overlap and also highlights CO3’s effectiveness in correcting the compositional limitations of the base model.
+Figure 4: Qualitative comparison of CO3 with competing methods on complex and rare concept prompts.
+Table 3: Ablations study conducted on the CO3 method. Starting with the base-model SDXL we progressively add different components.
+ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$
+Method A-A A-O O-O A-A A-O O-O
+SDXL 0.7820 1.5474 0.6789 0.6951 0.8654 0.4926 0.8435
++ Resampling 1.0881 1.6755 0.8263 0.7351 0.8666 0.4528 0.9437
++ Corrector 1.0630 1.6429 0.8949 0.7177 0.8794 0.4796 0.9463
++ weight-mod. 1.2342 1.6744 1.0158 0.7441 0.8878 0.5146 1.0118
+SDXL+Corrector 0.9543 1.5140 0.8894 0.7149 0.8801 0.4989 0.9086
++ weight-mod. 1.1732 1.5865 1.0067 0.7211 0.8701 0.5206 0.9797
+Figure 5: Qualitative results of CO3 with different configurations on SDXL.
+Qualitative Comparison: Figure 3 compares CO3 on Attend-Excite (simpler) prompts with several baselines.
 We illustrate common issues in compositional alignment—concept missing, attribute mixing, and object binding—across all categories.
-Divide-Bind and Attend-and-Excite enhance the attention layer of the base model, but often yield cropped or incomplete concepts (rows 1, 3).
-Composable Diffusion suffers from missing or merged concepts (rows 2, 3).
-SynGen and ToMe perform well on Object prompts of the form “a [color1] [object1] and a [color2] [object2],” but still exhibit overlapping or mixed concepts (rows 2, 3; columns 5, 6).
-In contrast, CO3 achieves stronger object binding in the Animals and Animal–Objects categories while also improving results on Objects.
-Figure 3: Qualitative comparison of different methods on simpler prompts.
-Figure 4 extends qualitative results to complex prompts from the Complex category of Huang et al.
-(2023), comparing CO3 with the top-performing SDXL models (Table 1).
-The first row shows prompts involving spatial relations (e.g., “on,” “to the left of”), while the second row depicts more descriptive prompts with multiple interacting concepts.
-SynGen and ToMe successfully bind concepts to their subjects but often miss concepts or relations.
-By mitigating dominant-concept bias, CO3 remains effective even for complex prompts with multiple relations.
-Figure 4: Qualitative comparison of CO3 with competing methods on complex prompts.
-Ablations Studies are preformed on the key components of CO3, starting from the base SDXL model (Table 2).
+Divide-Bind (Liu et al., 2023) and Attend-Excite (Chefer et al., 2023) enhance the attention layer of the base model, but often yield cropped or incomplete concepts (rows 1, 3).
+Comp-Diff (Liu et al., 2022) suffers from missing or merged concepts (rows 2, 3).
+SynGen (Rassin et al., 2023) and ToMe (taihang Hu et al., 2024) perform well on Object prompts of the form “a [color1] [object1] and a [color2] [object2],” but still exhibit overlapping or mixed concepts (rows 2, 3; columns 5, 6).
+In contrast, CO3 achieves stronger object binding in the Animals-Animals and Animals–Objects categories while also improving results on Objects-Objects.
+Figure 4 extends qualitative results to complex prompts from the Complex category of T2ICompbench and rare-concept prompts from Rarebench Park et al.
+(2024), comparing CO3 with the top-performing SDXL models (from Table 1).
+The left half shows prompts from T2ICompbench’s Complex category, which involves complex subject-attribute & inter-concept relations.
+Only CO3 is able to faithfully generate all concepts and relations, while the other methods often miss concepts or fail to capture relations.
+The right half shows prompts from RareBench, which involve rare or novel concepts.
+CO3 is able to generate extremely rare concepts such as ”banana shaped bottle” or ”Knighted turtle”, while the other methods often suffer from training data-bias – generating a regular banana or human knight instead.
+Ablations Studies are preformed on the key components of CO3, starting from the base SDXL model (Table 3 & Figure 5).
 Adding resampling in the first three DDIM steps substantially improves average performance, particularly for the Animals and Animal–Objects categories.
 Since image layout is largely determined in the early timesteps, CO3-resampler iteratively explores better initial noise configurations that align with the prompt.
-Incorporating CO3-corrector over the next seven steps further improves prompts requiring fine-grained details such as color or texture, boosting performance on the Objects category in both ImageReward and BLIP-VQA.
+Figure 6: Model Agnostic behavior: Qualitative comparison of generation from PixART-$\Sigma$(Chen et al., 2024) base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion (Comp-diff).
+Incorporating CO3-corrector over the next steps further improves prompts requiring fine-grained details such as color or texture, boosting performance on the Objects category in both ImageReward and BLIP-VQA.
+Table 4: Generalization of CO3 on PixArt and comparison with model-agnostic baselines. cDiff denotes Comp-diff (Liu et al., 2022).
+ImageReward $\uparrow$ BLIP-VQA $\uparrow$
+Method A-A A-O O-O A-A A-O O-O
+PixArt 1.066 1.697 1.023 0.691 0.861 0.559
+PixArt + cDiff 0.917 0.047 0.339 0.208 0.591 0.472
+PixArt + CO3 1.380 1.770 1.081 0.746 0.864 0.581
 Additional gains come from the Closeness-Aware Concept Weight Modulation strategy, which consistently benefits all categories.
 We hypothesize that as the diffusion trajectory evolves, the latent state often drifts closer to one concept mode $c_{i}$; the modulation mechanism assigns $c_{i}$ a more negative weight, repelling the latent and mitigating concept dominance (Tunanyan et al., 2023) commonly observed in stable-diffusion.
-BLIP-VQA ImageReward
-Methods animals animals-objects objects animals animals-objects objects Avg.
-SDXL 0.6951 0.8654 0.4926 0.7820 1.5474 0.6789 0.8435
-+ resampling 0.7351 0.8666 0.4528 1.0881 1.6755 0.8263 0.9407
-   + Corrector 0.7177 0.8794 0.4796 1.0630 1.6429 0.8949 0.9463
-   + weight-modulation 0.7441 0.8878 0.5146 1.2342 1.6744 1.0158 1.0118
-Table 2: Ablations study conducted on the CO3 method. Starting with the base-model SDXL we progressively add different components.
-Comparison with Optimization-Free Approaches: To further demonstrate CO3’s model-agnostic nature and generalization to other base models, we evaluate it on PixArt-$\Sigma$ (Chen et al., 2024).
-We compare against both the base model and the sampling-based, optimization-free Composable Diffusion (Liu et al., 2022).
-Unlike Composable Diffusion, which composes diffusion chains within the DDIM prediction step, CO3 applies composition as a corrector.
+Including the Corrector alone already improves performance in categories where attributes must attach to specific objects.
+This is most clearly visible in Rows 4–5 of the ablation table: prompts of the form “an [attribute1][object1] and [attribute2][object2]” show large gains, because early-timestep dominance (one object absorbing both attributes) is the failure mode that the Corrector resolves.
+Importantly, the average performance of SDXL + Corrector + weight-modulation (last row) is very close to SDXL + Resampling + Corrector + weight-modulation (Row 4), which indicates that the Corrector alone is powerful even without Resampling.
+However, the full improvement is achieved when the Corrector is applied after the Resampling steps.
+Comparison with Optimization-Free Approaches: To further demonstrate CO3’s model-agnostic nature and generalization to other base models, we evaluate it on PixArt-$\Sigma$ (Chen et al., 2024) (Table 4).
+We compare against both the base model and the sampling-based, optimization-free Composable Diffusion (Comp-diff) (Liu et al., 2022).
+Unlike Comp-diff, which composes diffusion chains within the DDIM prediction step, CO3 applies composition as a corrector.
 As shown in Fig.
-5, Composable Diffusion suffers from missing or mixing concepts, while CO3 consistently preserves all concepts with correct bindings in the output.
-Figure 5: Model Agnostic behavior: Qualitative comparison of generation from PixART-$\Sigma$(Chen et al., 2024) base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion.
+6, Comp-diff suffers from missing or mixing concepts, while CO3 consistently preserves all concepts with correct bindings in the output.
 5 Related Work
-Optimization-based correction approaches: (1) Text-embedding optimization works like T2I-Zero variants (Tunanyan et al., 2023), and style or prompt inversion (Li et al., 2023a), improve semantic alignment by optimizing or restructuring the text representation during inference.
+Optimization-based correction approaches: (1) Text-embedding optimization works like T2I-Zero variants (Tunanyan et al., 2023; Zhuang et al., 2024), and style or prompt inversion (Li et al., 2023a), improve semantic alignment by optimizing or restructuring the text representation during inference.
 These methods tweak or invert token embeddings so that the denoising trajectory better reflects entity–attribute bindings or multi-concept prompts.
@@
 InitNO (Guo et al., 2024) optimizes the initial noise so that sampling begins in more favorable regions that yield stronger, less-conflicted attention.
-Composable generation works: view conditional diffusion models as energy/score functions, enabling algebraic composition.
+Composable generation works: These works view conditional diffusion models as energy/score functions, enabling algebraic composition.
 Composed-Diffusion (Liu et al., 2022) frames score-composition under CFG and demonstrates test-time generalization, but suffers from concept mixing and missing.
 Follow-ups propose training-free and model-agnostic methods like energy-parameterized diffusion and Metropolis/MCMC-corrected samplers that markedly improve multi-condition generation (Du et al., 2024), but their performance is often poor (Chefer et al., 2023; Feng et al., 2022).
-Kwon & Ye (2025) uses a composition similar to our resampler but not in the context of a corrector.
+On a different line, while Skreta et al.
+(2025) proposes a density estimation approach for composing diffusion chains for concept interpolation, Park et al.
+(2024) interpolates distributions of frequent and rare-concepts for improved compositional generation.
+Kwon and Ye (2025) uses a composition similar to our resampler but not in the context of a corrector.
 Instead they use multiple DDIM forward-backward steps to sample initial noise.
@@
 Results demonstrate that CO3 outperforms baseline approaches for several diffusion model choices and serves as a plug-and-play module offering improved compositional generation performance.
-While our corrector’s closeness-aware weight modulation suppresses the concept mode overlap, the noise/score landscape is strongly biased by the training paradigms of diffusion models, i.e., the quantity and quality of the multi-concept bindings in the training set.
-This problem also manifests when presented with more unrealistic prompts, resulting in poor text/concept alignment.
-As shown in Fig.
-7, while our work offers a significant boost to multi-concept compositional generation, clearly, there are failure cases that are not yet addressed.
-However, we believe that more advanced energy-based compositional samplers—that explicitly take into account the probabilities of the landscape during sampling—can further help in avoiding such cases.
-We hope to investigate these special cases in the future.
-We also hope to explore applications beyond composition using the CO3 corrector.
-7 Reproducibility Statement
+While our corrector suppresses concept overlap, the model’s predicted score landscape remains biased by training data quality, occasionally leading to poor alignment in unrealistic prompts (Fig.
+8).
+While our work offers a significant boost to multi-concept compositional generation, failure cases still persist and we believe advanced energy-based compositional samplers that explicitly model landscape probabilities can further mitigate these issues.
+We intend to investigate these edge cases and explore applications beyond compositional generation in future work.
+Acknowledgments
+Acknowledgments: This work was partially supported by NSF #2008338, #1909568, #2148583, and #MRI-2018966.
+This work used DELTA at NCSA through allocation CIS230230 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S.
+National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296
+Reproducibility Statement
 All method implementations, inference pipelines are included in the supplementary material.
@@
 Lemma 1.
-Let’s define CFG as a function at time $t$ as, $f_{CFG}({\epsilon};\lambda)=$ ${\epsilon}^{\phi}+\lambda({\epsilon}-{\epsilon}^{\phi})$.
-Then for any $\lambda$ and $K>1$, $\sum_{k}{w_{k}f_{CFG}}=f_{CFG}(\sum_{k}{w_{k}{\epsilon}^{k}})$ only if $\sum_{k}w_{k}=1$.
+Let’s define CFG as a function at time $t$ as, $f_{CFG}({\epsilon}_{t};\lambda)=$ ${\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}-{\epsilon}_{t}^{\phi})$.
+Then for any $\lambda$ and $K>1$, $\sum_{k}{w_{k}f_{CFG}({\epsilon}_{t}^{k};\lambda)}=f_{CFG}(\sum_{k}{w_{k}{\epsilon}_{t}^{k}};\lambda)$ only if $\sum_{k}w_{k}=1$.
 Proof.
-Let, $\tilde{{\epsilon}}^{\lambda}_{t}=\sum_{k}{w_{k}f_{CFG}}$. Then,
-$\tilde{{\epsilon}}^{\lambda}_{t}$ $=\sum_{k}w_{k}({\epsilon}^{\phi}+\lambda({\epsilon}^{k}-{\epsilon}^{\phi}))$
-$=\sum_{k}w_{k}{\epsilon}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-\sum_{k}w_{k}{\epsilon}^{\phi})$
-$={\epsilon}_{\phi}\sum_{k}w_{k}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-{\epsilon}^{\phi}\sum_{k}w_{k})$
-$={\epsilon}_{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}^{k}-{\epsilon}^{\phi})\quad\text{if }\sum_{k}w_{k}=1$
-$=f_{CFG}(\sum_{k}w_{k}{\epsilon}^{k})\quad\text{if }\sum_{k}w_{k}=1$
-If $\sum_{k}w_{k}\neq 1$, then $\tilde{{\epsilon}}^{\lambda}_{t}\neq f_{CFG}(\sum_{k}w_{k}{\epsilon}^{k})$. ∎
+Let, $\tilde{{\epsilon}}^{\lambda}_{t}=\sum_{k}{w_{k}f_{CFG}({\epsilon}_{t}^{k};\lambda)}$. Then,
+$\tilde{{\epsilon}}^{\lambda}_{t}$ $=\sum_{k}w_{k}({\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{k}-{\epsilon}_{t}^{\phi}))$
+$=\sum_{k}w_{k}{\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{k}-\sum_{k}w_{k}{\epsilon}_{t}^{\phi})$
+$={\epsilon}_{t}^{\phi}\sum_{k}w_{k}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{k}-{\epsilon}_{t}^{\phi}\sum_{k}w_{k})$
+$={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{k}-{\epsilon}_{t}^{\phi})\quad\text{if }\sum_{k}w_{k}=1$
+$=f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{k};\lambda)\quad\text{if }\sum_{k}w_{k}=1$
+If $\sum_{k}w_{k}\neq 1$, then $\tilde{{\epsilon}}^{\lambda}_{t}\neq f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{k};\lambda)$. ∎
 Remark: One immediate implication of this lemma is that, whenever we sample from a composed distribution with the CFG composed score of the form $\tilde{{\epsilon}}^{\lambda}=w_{1}{\epsilon}^{\lambda,1}+w_{2}{\epsilon}^{\lambda,2}+\dots+w_{K}{\epsilon}^{\lambda,K}$, then $\tilde{{\epsilon}}^{\lambda}$ corresponds to the CFG noise for the distribution $p(x|C)=\prod_{i}p(x\mid c_{i})^{w_{i}}$ only when $\sum_{i}w_{i}=1$.
-Lemma 3.
-Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$.
+Proposition 2.
+(Restated) Let $\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c}]:=x_{t}-\sigma_{t}\ {\epsilon}_{t}^{\lambda,c}$ be the tweedie mean from CFG composed noise $\tilde{{\epsilon}_{t}}^{\lambda}={\epsilon}_{t}^{\phi}+\lambda({\epsilon}_{t}^{C}-{\epsilon}_{t}^{\phi})$ for some $\lambda$.
 Let, $\tilde{\hat{x}}_{tweedie}$ be the composed tweedie-mean defined as $\tilde{\hat{x}}_{tweedie}=\sum_{k}w_{k}\hat{x}_{tweedie}[{\epsilon}_{t}^{\lambda,c_{k}}]$.
@@
   2. b)
-    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
+    CO3-resampler: $\tilde{\hat{x}}_{tweedie}=-\lambda\ \sigma_{t}\ \sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}$ is weighted noise if and only if $\sum_{k}w_{k}=0$.
 Proof.
@@
 […]
-$=x_{t}-\sigma_{t}\ f_{CFG}(\sum_{k}w_{k}{\epsilon}^{c_{k}})\quad\text{if }\sum_{k}w_{k}=1$
+$=x_{t}-\sigma_{t}\ f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}})\quad\text{if }\sum_{k}w_{k}=1$
 So, we have $\tilde{\hat{x}}_{tweedie}=x_{t}-\sigma_{t}\ \tilde{{\epsilon}_{t}}^{\lambda,C}$ where $\tilde{{\epsilon}_{t}}^{\lambda,C}=f_{CFG}(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}})={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$.
@@
 […]
-A.3 More Implementation Details
-A.3.1 CO3 Implementation Details
+A.3 Comparison of compositions in score space
+In this section we compare our composition framework with Composable-Diffusion. As already described in section 3, we propose composition in Tweedie-denoised space as
+$\tilde{x}_{tweedie}=w_{0}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,C}]\;+\;w_{1}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{1}}]\;+\;\dots\;+\;w_{K}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{K}}]$ (20)
+which leads to
+$\tilde{x}_{tweedie}$ $=x-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}\text{ iff }\sum w_{k=0}^{K}=1$ (22)
+$\text{ where }\tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ $={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$ (23)
+Thus, when mixing scores from different conditional distributions, the above condition need to be satisfied for $\tilde{x}_{tweedie}$ to correspond to any coherent joint conditional model.
+Contrast this with the Composable-Diffusion’s (Liu et al.
+(2022)) noise/score composition:
+$\tilde{\epsilon}^{\lambda,C}_{t,compdiff}\;=\;\epsilon_{t}^{\phi}+\lambda_{1}\bigl(\epsilon_{t}^{c_{1}}-\epsilon_{t}^{\phi}\bigr)+\lambda_{2}\bigl(\epsilon_{t}^{c_{2}}-\epsilon_{t}^{\phi}\bigr)+\dots+\lambda_{K}\bigl(\epsilon_{t}^{c_{K}}-\epsilon_{t}^{\phi}\bigr)$ (24)
+The above composition allows arbitrary weights, which in general violate the unit-sum constraint.
+In such cases, the result is simply a heuristic linear combination of scores, and the resultant Tweedie mean isn’t valid and doesn’t correspond to any well-defined distribution under the given forward process.
+By contrast, CO3 composes Tweedie-means from different conditional distributions.
+Lemma 1 shows that these weights have to satisfy the same unit-sum condition as above for the resultant composed Tweedie-mean to be a valid one.
+CO3 maintains this condition and this preserves the update on the Tweedie manifold of the underlying sampler.
+A.4 More Implementation Details
+A.4.1 CO3 Implementation Details
 For the results in Table 1 we use $T_{c}=10$(number of time-steps to correct), $T_{r}=3$(number of resampling steps) and P=5P=5(number of corrector iteration).
@@
 For example, if $C$ is ”a black cat and a brown dog”, we consider $c_{1}=$”cat” and $c_{2}=$”dog”.
-A.3.2 Baseline Methods
+A.4.2 Baseline Methods
 Attend-Excite (Chefer et al., 2023), Divide-Bind (Liu et al., 2023), InitNo (Guo et al., 2024), SynGen (Rassin et al., 2023) and Composable Diffusion (Liu et al., 2022) are methods based on SD1.5.
 We run experiments on these models using their publicly available code.
-ToMe (taihang Hu et al., 2024) is SDXL based we use their publicly available code-base implementation.
+Tweediemix (Kwon and Ye, 2025), Magnet Zhuang et al.
+(2024), R2F (Park et al., 2024) and ToMe (taihang Hu et al., 2024) are SDXL based we use their publicly available code-base implementation.
 We also adapted Composable Diffusion and SynGen to SDXL.
-A.4 Time Complexity Analysis
-We conduct an experiment to analyze the time-complexity of each SDXL based methods.
-The Table 3 indicates the average time (in secs) taken by the model per sample.
-The experiments are conducted on NVIDIA A100 gpu.
-Method Inference Steps Time Cost (sec) Animals Animals-Objects Objects
-SDXL 50 7.20 0.6950 0.8654 0.4926
-Composable Diffusion 50 9.80 0.2846 0.5656 0.4529
-SynGen(SDXL) 50 11.48 0.6816 0.8578 0.4652
-ToMe 50 16.58 0.6257 0.8808 0.6440
-CO3(ours) 50 19.9 0.7441 0.8878 0.5146
-Table 3: Comparison of different methods across efficiency and compositionality metrics.
-A.5 Comparison of compositions in score space
-In this section we compare our composition framework with Composable-Diffusion. As already described in section 3, we propose composition in Tweedie-denoised space as
-$\tilde{x}_{tweedie}=w_{0}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,C}]\;+\;w_{1}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{1}}]\;+\;\dots\;+\;w_{K}\,\hat{x}_{tweedie}[\epsilon_{t}^{\lambda,c_{K}}]$ (20)
-which leads to
-$\tilde{x}_{tweedie}$ $=x-\sigma_{t}\ \tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}\text{ iff }\sum w_{k=0}^{K}=1$ (22)
-$\text{ where }\tilde{{\epsilon}}_{t}^{\tilde{\lambda},C}$ $={\epsilon}_{t}^{\phi}+\lambda(\sum_{k}w_{k}{\epsilon}_{t}^{c_{k}}-{\epsilon}_{t}^{\phi})$ (23)
-Contrast this with the Composable-Diffusion’s noise/score composition:
-$\tilde{\epsilon}^{\lambda,C}_{t,compdiff}\;=\;\epsilon_{t}^{\phi}+\lambda_{1}\bigl(\epsilon_{t}^{c_{1}}-\epsilon_{t}^{\phi}\bigr)+\lambda_{2}\bigl(\epsilon_{t}^{c_{2}}-\epsilon_{t}^{\phi}\bigr)+\dots+\lambda_{K}\bigl(\epsilon_{t}^{c_{K}}-\epsilon_{t}^{\phi}\bigr)$ (24)
-For arbitrary weights $\lambda_{k}$}, this cannot be expressed in the from equation 23. This proves that $\tilde{\epsilon}^{\lambda,C}_{t,compdiff}$ doesn’t lead to a valid Tweedie-mean.
-A.6 More Visualizations
-Figure 6: Qualitative results of SDXL base diffusion model, and SDXL + CO3 for animal-animal,animal-object, and object-object prompt categories
-A.6.1 CO3 Failure Cases
-Despite correcting for the “problematic” modes, there still remain open challenges in image composition as shown in Figure 7.
-The score and, hence, the correction landscape is heavily influenced by the training schemes employed in diffusion model training, i.e., the quantity and quality of the multi-concept bindings in the training set.
-In addition, the usage of unrealistic prompts perhaps not encountered in training also results in poor text/concept alignment.
-We leave this investigation for the future.
-Figure 7: Failure scenarios of PixART-$\Sigma$ base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion.
-A.7 Ablations on Hyperparameters
+Table 5: Human Evaluation results on Attend-Excite prompts. Higher the score, the better.
+Method Animals Animals- Objects Objects
+SDXL 0.5266 0.6050 0.7083
+SynGen 0.3667 0.4366 0.7183
+Magnet 0.3700 0.5683 0.6849
+ToMe 0.4383 0.6083 0.6166
+CO3(ours) 0.8817 0.7466 0.8250
+A.4.3 Additional Human Evaluation
+We conducted new human-evaluation studies (with 11 participants) across all benchmarks considered in this paper.
+For evaluation, we follow the protocol from R2F Park et al.
+(2024).
+For each prompt, images from CO3 and baseline methods are displayed side by side to enable direct comparison.
+Method names are anonymized and their order is randomly shuffled to reduce bias.
+Participants provide ratings in $\{0,1,2,3,4,5\}$, which are then normalized to $[0,1]$.
+Results are shown below (higher is better).
+Results on T2ICompbench and RareBench are included in Table 2 of the main text.
+Table 5 shows results on Attend-Excite prompts.
+A.5 Generalization to higher-order solvers and low inference steps regimes
+In addition to the main 50-step SDXL with DDIM experiments, we include results using both (i) low-step DDIM and (ii) low-step high-order DPM++ solvers (10–20 steps).
+Please note that, since DDIM sampling has direct connections with denoising and generation via the Tweedie mean (Section 2 in the main paper), CO3 applies naturally to DDIM.
+However, our experiments below show that CO3 continues to provide consistent improvements under higher-order solvers and aggressive sampling budgets as shown in Table 6 & Table 7.
+Table 6: Performance of CO3 with DPM++ solver and in the low inference steps regime.
+Steps ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$
+Method Animals Animals Objects Objects Animals Animals Objects Objects
+SD-DDIM 50 0.7820 1.5574 0.6789 0.6950 0.8658 0.4925 0.8452
+CO3-DDIM 50 1.2341 1.6743 1.0158 0.7441 0.8878 0.5146 1.0112
+SD-DPM++ 20 0.7089 1.5652 0.6757 0.6974 0.8537 0.4984 0.8332
+CO3-DPM++ 20 1.1424 1.5249 0.9537 0.7382 0.8842 0.4944 0.9563
+SD-DPM++ 10 0.6834 1.5074 0.6681 0.6768 0.8537 0.4974 0.8145
+CO3-DPM++ 10 1.0371 1.4520 0.9206 0.7331 0.8765 0.4443 0.9106
+Table 7: Performance of CO3 with DDIM solver in the low inference steps regime.
+ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$
+Method Steps Animals Animals Objects Objects Animals Animals Objects Objects
+SD w/ DDIM 50 0.7820 1.5574 0.6789 0.6950 0.8658 0.4926 0.8452
+CO3 w/ DDIM 50 1.2341 1.6743 1.0158 0.7441 0.8878 0.5146 1.0112
+SD w/ DDIM 20 0.6965 1.4567 0.5516 0.6832 0.8620 0.4947 0.7907
+CO3 w/ DDIM 20 0.8532 1.6440 0.7481 0.6698 0.8611 0.4315 0.8671
+SD w/ DDIM 10 0.4906 1.1503 0.1131 0.6428 0.8378 0.4752 0.6183
+CO3 w/ DDIM 10 0.5872 1.5900 0.7117 0.6294 0.8589 0.4453 0.8037
+A.6 PROMPTS WITH AND WITHOUT BACKGROUNDS
+Figure 7: Concept obedient behavior of CO3: CO3 tends to strongly obey the concepts specified in the prompt; when the background is not specified it is conservative with hallucinating the background.
+CO3 is strongly concept obedient.
+Many CO3 generated images from Attend-Excite prompts seem to lack background details.
+To investigate this behavior, we added background concepts to the same prompts shown in Figure 6 of the main text.
+Figure 7 shows the resulated images from this experiment.
+We observe that, CO3 tends to strongly obey the concepts specified in the prompt; when the background is not specified it is conservative with hallucinating the background.
+However, if the background is explicitly specified, CO3 brings the same degree of richness and quality to both the foreground and the background.
+A.7 Time and Memory Complexity
+Table 8: Runtime and memory comparison between SDXL, ToMe, and CO3 at different sampling budgets.
+ImageReward $\uparrow$ BLIP-VQA $\uparrow$ Avg $\uparrow$ Time(s) $\downarrow$ Memory $\downarrow$
+Method Animals Animals Objects Objects Animals Animals Objects Objects
+SDXL (50 steps) 0.4906 1.1503 0.1131 0.6428 0.8378 0.4752 0.6183 7.20 8.5GB
+ToMe (50 steps) 0.3894 1.5736 1.0117 0.6257 0.8807 0.6440 0.8542 16.58 18GB
+CO3 (50 steps) 1.2341 1.6743 1.0158 0.7440 0.8877 0.5145 1.0120 19.90 9.5GB
+CO3 (20 steps) 0.8532 1.6440 0.7481 0.6698 0.8611 0.4315 0.8679 10.65 9.5GB
+We compare CO3 with its most competitive baseline ToMe taihang Hu et al.
+(2024) in terms of runtime peak memory and time requirement to generate one sample in Table 8.
+The experiments are conducted on a NVIDIA A100 gpu.
+Although CO3 takes more time compared to ToMe with 50 time-steps, it requires significantly less memory.
+CO3 is lightweight, and its peak memory usage is only slightly above that of the base SDXL model.
+ToMe, being a gradient-based method, has substantial memory requirements.
+We also emphasize that CO3 is model-agnostic and can be applied to any diffusion-based model.
+In contrast, ToMe is model-architecture dependent, as its attribute-binding mechanism operates by manipulating internal attention activations of the base model.
+For a fair comparison, we additionally tested CO3 with 20 inference steps.
+As shown in the table below, the average performance of CO3 with 20 steps remains comparable to, and in many cases better than, ToMe with 50 steps.
+CO3 with $20$ steps takes $10.65$ seconds, which is faster than ToMe.
+A.8 Ablations on Hyperparameters
 In this section, we analyze the contribution of the following five factors to the performance of our CO3 corrector.
 Notation recap: $\beta$ is the exponential decay factor in the affinity scores; $\lambda$ scales the composed_score for CFG; num_resampling is the number of resampling steps at the start of diffusion; num_ts is the number of early timesteps where the corrector is used; num_steps is the number of iterations per corrector application.
-A.7.1 Number of resampling steps
+A.8.1 Number of resampling steps
 Table 9 shows results for different resampling steps.
@@
 Frozen settings num_resampling ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects $\beta=0.8$, num_ts=7, $\lambda=0.8$, num_steps=5 3 1.2218 1.7020 0.9405 0.7375 0.8811 0.4782 4 1.1705 1.7031 0.9512 0.7259 0.8828 0.4696 $\beta=0.3$, num_ts=7, $\lambda=0.8$, num_steps=5 3 1.1484 1.6915 0.9005 0.7240 0.8810 0.4691 4 1.1297 1.6948 0.8996 0.7145 0.8774 0.4697 $\beta=1.1$, num_ts=7, $\lambda=0.8$, num_steps=5 3 1.2566 1.7062 0.9916 0.7422 0.8813 0.4834 4 1.1927 1.7164 0.9934 0.7292 0.8835 0.4811
-A.7.2 Corrector application timesteps
+A.8.2 Corrector application timesteps
 In CO3, the corrector is applied during the first num_ts diffusion steps.
@@
 num_ts ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects 4 1.1862 1.7136 0.8885 0.7430 0.8783 0.4630 6 1.2510 1.7117 0.9750 0.7432 0.8801 0.4812 7 1.2566 1.7062 0.9916 0.7422 0.8813 0.4834 10 1.3149 1.6816 1.0095 0.7576 0.8832 0.5074
-A.7.3 Corrector iterations
+A.8.3 Corrector iterations
 Table 11 reports the effect of increasing the number of corrector iterations.
@@
 Fixed settings num_steps ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects (a) $\beta=1.1$, num_ts=4, num_resampling=2, $\lambda=0.9$ 7 1.2622 1.7129 0.9959 0.7542 0.8843 0.4707 10 1.2452 1.6994 0.9753 0.7448 0.8809 0.4706 15 1.2151 1.6701 0.9437 0.7318 0.8781 0.4556 (b) $\beta=1.1$, num_ts=4, num_resampling=2, $\lambda=0.8$ 7 1.2755 1.7049 0.9424 0.7538 0.8762 0.4611 10 1.2503 1.6969 0.9191 0.7446 0.8798 0.4617 15 1.2480 1.6658 0.9063 0.7356 0.8750 0.4577
-A.7.4 Exponential decay factor $\beta$ of affinity scores
+A.8.4 Exponential decay factor $\beta$ of affinity scores
 We vary $\beta$ in  equation 14 from 0.3 to 1.1.
-Tables 7 and 8 demonstrate that the performance generally increases with larger $\beta$, with the strongest gains in the animals and objects categories, while a few settings exhibit minor regressions.
+Tables 12 and 13 demonstrate that the performance generally increases with larger $\beta$, with the strongest gains in the animals and objects categories, while a few settings exhibit minor regressions.
 Table 12: Exponential decay factor $\beta$ with all other settings fixed: num_resampling=3, num_steps=5, $\lambda=0.8$.
@@
 num_ts $\beta$ ImageReward (↑) BLIP-VQA (↑) Animals Animals_Objects Objects Animals Animals_Objects Objects 6 0.3 1.1437 1.7020 0.9518 0.7248 0.8816 0.4764 0.7 1.2051 1.7130 0.9860 0.7407 0.8851 0.4788 0.8 1.2037 1.7110 0.9916 0.7388 0.8832 0.4894 0.9 1.2003 1.7159 0.9883 0.7416 0.8842 0.4891 1.1 1.2098 1.7180 0.9956 0.7371 0.8841 0.4855 7 0.3 1.1273 1.6938 0.9490 0.7274 0.8848 0.4796 0.8 1.2142 1.7000 0.9937 0.7363 0.8845 0.4952 1.1 1.2331 1.7046 1.0148 0.7496 0.8864 0.4945
+A.8.5 CO3 Failure Cases
+Despite correcting for the “problematic” modes, there still remain open challenges in image composition as shown in Figure 8.
+The score and, hence, the correction landscape is heavily influenced by the training schemes employed in diffusion model training, i.e., the quantity and quality of the multi-concept bindings in the training set.
+In addition, the usage of unrealistic prompts perhaps not encountered in training also results in poor text/concept alignment.
+We leave this investigation for the future.
+Figure 8: Failure scenarios of PixART-$\Sigma$ base diffusion model, PixART-$\Sigma$ + CO3, and PixART-$\Sigma$ + Composable Diffusion.
+A.9 More Qualitative results on Multi-concept Prompts
+Figure 9: Qualitative results of SDXL base diffusion model, and SDXL + CO3 for animal-animal,animal-object, and object-object categories from Attend-Excite prompts.
+Figure 10: Qualitative results on Complex category from T2Icompbench
+Figure 11: Qualitative results on RareBench prompts.
+We added more qualitative comparison results on the three categories from Attend-Excite prompts in Figure 9, Complex category from T2Icompnench in Figure 10.
+Additional results on Rarebench prompts are included in Figure 11.